MicroEMACS

Full Screen Text Editor Reference Manual - Preliminary Version Version 3.12 April 22, 1993

Copyright \begin{matrix}c \\ ◯ \end{matrix} 1988 - 1993 by Daniel M. Lawrence

Reference Manual \begin{matrix}c \\ ◯ \end{matrix} 1988 - 1993

by Brian Straight and Daniel M. Lawrence All Rights Reserved

\begin{matrix}c \\ ◯ \end{matrix} 1988 - 1993 by Daniel M. Lawrence MicroEMACS 3.12 can be copied and distributed freely for any non-commercial purposes. Commercial users may use MicroEMACS 3.12 inhouse. Shareware distributors may redistribute MicroEMACS 3.12 for media costs only. MicroEMACS 3.12 can only be incorporated into commercial software or resold with the permission of the current author.

MicroEMACS Reference Manual 2 MicroEMACS Reference Manual 2

CONTENTS  MicroEMACS Reference Manual

Contents

1 Installation 1 1.1 MSDOS - IBM-PCs . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 1 1.2 UNIX . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 1 1.3 ATARI ST . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 1

2 Basic Concepts 3 2.1 Keys and the Keyboard . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 3 2.2 Getting Started . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 3 2.3 Parts and Pieces . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 3 2.4 Entering Text . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 4 2.5 Basic cursor movement . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 5 2.6 Saving your text . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 5

3 Basic Editing–Simple Insertions and Deletions 7 3.1 A Word About Windows, Buffers, Screens, and Modes . . . . . . . . . . . . . . . . . . . . . . . . . . . . 7 3.2 Insertions . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 7 3.3 Deletions . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 8

4 Using Regions 11 4.1 Defining and Deleting a Region . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 11 4.2 Yanking a Region . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 12

5 Search and Replace 13 5.1 Forward Search . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 13 5.2 Exact Searches . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 14 5.3 Backward Search . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 14 5.4 Searching and Replacing . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 14 5.5 Query-Replace . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 14

6 Windows 17 6.1 Creating Windows . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 17 6.2 Deleting Windows . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 18 6.3 Resizing Windows . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 18 6.4 Repositioning within a Window . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 18

MicroEMACS Reference Manual  CONTENTS

7 Using a Mouse 21 7.1 Moving around with the mouse . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 21 7.2 Dragging around . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 21 7.3 Cut and Paste . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 22 7.4 Screens . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 22 7.5 Resizing a Screen . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 22 7.6 Moving a Screen . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 22 7.7 Creating a Screen . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 23 7.8 Switching to a Screen . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 23 7.9 Deleting a Screen . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 23 8 Buffers 25

9 Modes 27 9.1 ASAVE mode . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 27 9.2 CMODE mode . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 27 9.3 CRYPT mode . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 28 9.4 EXACT mode . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 28 9.5 MAGIC mode . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 28 9.6 OVER mode . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 29 9.7 WRAP mode . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 29 9.8 VIEW mode . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 30

10 Files 31

11 Screen Formatting 33 11.1 Wrapping Text . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 33 11.2 Reformatting Paragraphs . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 33 11.3 Changing Case . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 33 11.4 Tabs . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 34

12 Access to the Outside World 37

13 Keyboard Macroes 39

14 MicroEMACS Procedures 41 14.1 Constants . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 41 14.2 Variables . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 42

14.2.1 Environmental Variables . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 42 14.2.2 User variables . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 45 14.2.3 Buffer Variables . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 46 14.2.4 Interactive variables . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 46

14.3 Functions . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 47 14.4 Directives . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 49

14.4.1 !ENDM Directive . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 49 14.4.2 !FORCE Directive . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 49 14.4.3 !IF, !ELSE, and !ENDIF Directives . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 50 14.4.4 !GOTO Directive . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 50 14.4.5 !WHILE and !ENDWHILE Directives . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 51 14.4.6 !BREAK Directive . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 51 14.4.7 !RETURN Directive . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 51

15 Debugging MicroEMACS Procedures 53

16 Key Bindings, What they are and why 55

A MicroEMACS Command Line Switches and Startup Files 57

B Command Completion 59

C MicroEMACS Commands 61

D MicroEMACS Bindings 67

E Numeric Arguments to Commands 71 E.1 Commands unaffected by numeric arguments. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 74

F Supported machines 75 G Function Keys 77

H Machine Dependent Notes 79 H.1 IBM-PC/XT/AT and its clones . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 79 H.2 HP 150 . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 81 H.3 Atari 520/1040ST . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 82 H.4 Amiga 1000 . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 83 H.5 UNIX V5, V7, and BSD4.[23] . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 84 H.6 DEC VMS operating system . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 85

I Mode Flags 91

Introduction

MicroEMACS is a tool for creating and changing documents, programs, and other text files. It is both relatively easy for the novice to use, but also very powerful in the hands of an expert. MicroEMACS can be extensively customized for the needs of the individual user.

MicroEMACS allows several files to be edited at the same time. The screen can be split into different windows and screens, and text may be moved freely from one window on any screen to the next. Depending on the type of file being edited, MicroEMACS can change how it behaves to make editing simple. Editing standard text files, program files and word processing documents are all possible at the same time.

There are extensive capabilities to make word processing and editing easier. These include commands for string searching and replacing, paragraph reformatting and deleting, automatic word wrapping, word move and deletes, easy case controlling, and automatic word counts.

For complex and repetitive editing tasks editing macroes can be written. These macroes allow the user a great degree of flexibility in determining how MicroEMACS behaves. Also, any and all the commands can be used by any keystroke by changing, or rebinding, what commands various keys invoke.

Special features are also available to perform a diverse set of operations such as file encryption, automatic backup file generation, entabbing and detabbing lines, executing operating system commands and filtering of text through other programs (like SORT to allow sorting text).

History

EMACS was originally a text editor written by Richard Stallman at MIT in the early 1970s for Digital Equipment computers. Various versions, rewrites and clones have made an appearance since.

This version of MicroEMACS is derived from code written by Dave G. Conroy in 1985. Later modifications were performed by Steve Wilhite and George Jones. In December of 1985 Daniel Lawrence picked up the then current source (version 2.0) and made extensive modifications and additions to it over the course of the next eight years. Updates and support for the current version are still available. Commercial support and usage and resale licences are also available. The current program author can be contacted by writing to:

              USMAIL: Daniel Lawrence
                       617 New York St
                       Lafayette, IN 47901
              UUCP:   pur-ee!mdbs!dan
              ARPA:   mdbs!dan@ee.ecn.purdue.edu

Support is provided through:

              The Programmer's Room
              Opus 201/10
              300/1200/2400 and 9600 (Hayes V series only)
              (317) 742-5533  no parity  8 databits  no stop bits

Credits

Many people have been involved in creating this software and we wish to credit some of them here. Dave Conroy, of course, wrote the very first version of MicroEMACS, and it is a credit to his clean coding that so much work was able to be done to expand it. John Gamble is responsible for writing the MAGIC mode search routines, and for maintaining all the search code. Jeff Lomicka wrote the appendix on DEC VMS and has supplied a lot of code to support VMS and the ATARI 1040ST versions. Curtis Smith wrote the original VMS code and help support the Commodore AMIGA. Also Lance Jones has done a lot of work on the AMIGA code. Professor Suresh Konda at Carnegie Mellon University has put a lot of effort into writing complex macroes and finding all the bugs in the macro language before anyone else does.

A special thanks to Dana Hoggatt who has provided an almost daily sounding board for ideas, algorythms and code. He is responsible for the encryption code directly and has prodded me into adding many features with simple but poignant questions (Dan? How do we move the upper left corner of the screen? … which forced me to write the text windowing system).

Pierre Perrot dealt with my restrictive path to a generalized windowing version, and produced an excellent version for MicroSoft Windows. He continues to assist with this environment, forcing me to keep pace with him, making all the version more interesting.

As to people sending source code and text translations over computer networks like USENET and ARPA net, there are simply more than can be listed here. [The comments in the edit history in the history.c file mention each and the piece they contributed]. All these people should be thanked for the hard work they have put into MicroEMACS.

Daniel M. Lawrence 8

Installation  MicroEMACS Reference Manual

Chapter 1

Installation

MicroEMACS is a programmer’s text editor which is very powerfull, customizable, and exists for a large number of different types of computer systems. It is particularly usefull for people who work on a lot of different computers and want to have a familiar and powerful editor which works identically no matter what computer they are using. But before using MicroEMACS, you must INSTALL it on your computer system. Since each computer is different, there is usually a different way to install MicroEMACS for each type of computer.

1.1 MSDOS - IBM-PCs

1.2 UNIX

1.3 ATARI ST

MicroEMACS Reference Manual  Installation

2

Basic Concepts  MicroEMACS Reference Manual

Chapter 2

Basic Concepts

The current version of MicroEMACS is 3.12 (Third major re-write, twelveth public release), and for the rest of this document, we shall simply refer to this version as ”EMACS”. Any modifications for later versions will be in the file README on the MicroEMACS distribution disk.

2.1 Keys and the Keyboard

Many times throughout this manual we will be talking about commands and the keys on the keyboard needed to use them. There are a number of ”special” keys which can be used and are listed here:

<NL> NewLine which is also called RETURN, ENTER, or <NL>, this key is used to end different commands.

^The control key can be used before any alphabetic character and some symbols. For example, ^C means to hold down the <CONTROL> key and type the C key at the same time.

^X The CONTROL-X key is used at the beginning of many different commands.

META or M- This is a special EMACS key used to begin many commands. This key is pressed and then released before typing the next character. On most systems, this is the <ESC> key, but it can be changed. (consult appendix E to learn what key is used for META on your computer).

Whenever a command is described, the manual will list the actual keystrokes needed to execute it in boldface using the above conventions, and also the name of the command in italics.

2.2 Getting Started

In order to use EMACS, you must call it up from your system or computer’s command prompt. On UNIX and MSDOS machines, just type ”emacs” from the command prompt and follow it with the <RETURN> or <ENTER> key (we will refer to this key as <NL> for ”new-line” for the remainder of this manual). On the Macintosh, the Amiga, the ATARI ST, and under OS/2 and other icon based operating systems, double click on the uEMACS icon. Shortly after this, a screen similar to the one below should appear.

2.3 Parts and Pieces

The screen is divided into a number of areas or windows. On some systems the top window contains a function list of unshifted and shifted function keys. We will discuss these keys later. Below them is an EMACS mode line which, as we

MicroEMACS Reference Manual  Basic Concepts

===============================================================================
f1 search-> f2 <-search |     MicroEMACS:  Text Editor
f3 hunt->   f4 <-hunt   |
f5 fkeys     f6 help      |  Available function key Pages include:
f7 nxt wind f8 pg[     ] |     WORD  BOX  EMACS  PASCAL  C  Cobol  Lisp
f9 save      f10 exit     |  [use the f8 key to load Pages]
===============================================================================
   MicroEMACS 3.12 ()       Function Keys
===============================================================================

===============================================================================
---- MicroEMACS 3.12 () -- Main -----------------------------------------------
===============================================================================

Figure 1: EMACS screen on an IBM-PC

will see, informs you of the present mode of operation of the editor — for example ”(WRAP)” if you set EMACS to wrap at the end of each line. Under the mode line is the text window where text appears and is manipulated. Since each window has its own mode line, below the text window is it’s mode line. The last line of the screen is the command line where EMACS takes commands and reports on what it is doing.

2.4 Entering Text

Entering text in EMACS is simple. Type the following sentence fragment:

Fang Rock lighthouse, center of a series of mysterious and

The text is displayed at the top of the text window. Now type:

terrifying events at the turn of the century

Notice that some of your text has disappeared off the left side of the screen. Don’t panic — your text is safe!!! You’ve just discovered that EMACS doesn’t ”wrap” text to the next line like most word processors unless you hit <NL>. But since EMACS is used for both word processing, and text editing, it has a bit of a dual personality. You can change the way it works by setting various modes. In this case, you need to set WRAP mode, using the add-mode command, by typing ^XM. The command line at the base of the screen will prompt you for the mode you wish to add. Type wrap followed by the <NL> key and any text you now enter will be wrapped. However, the command doesn’t wrap text already entered. To get rid of the long line, press and hold down the <BACKSPACE> key until the line is gone. Now type in the words you deleted, watch how EMACS goes down to the next line at the right time. (In some versions of EMACS, WRAP is a default mode in which case you don’t have to worry about the instructions relating to adding this mode.)

Now let’s type a longer insert. Hit <NL> a couple of times to tab down from the text you just entered. Now type the following paragraphs. Press <NL> twice to indicate a paragraph break.

Fang Rock lighthouse, center of a series of mysterious and terrifying events at the turn of the century, is built on a rocky island a few miles of the Channel coast. So small is the island that wherever you stand its rocks are wet with sea spray. The lighthouse tower is in the center of the island. A steep flight of steps leads to the heavy door in its base. Winding stairs lead up to the crew room.

2.5 Basic cursor movement

Now let’s practice moving around in this text. To move the cursor back to the word ”Winding,” enter M-B previous-word. This command moves the cursor backwards by one word at a time. Note you have to press the key combination every time the cursor steps back by one word. Continuously pressing META and toggling B produces an error message. To move forward to the word ”stairs” enter M-F next-word, which moves the cursor forward by one word at a time.

Notice that EMACS commands are usually mnemonic — F for forward, B for backward, for example.

To move the cursor up one line, enter ^P previous-line , down one line ^N next-line . Practice this movement by moving the cursor to the word ”terrifying” in the second line.

The cursor may also be moved forward or backward in smaller increments. To move forward by one character, enter ^F forward-character , to move backward, ^B backward-character . EMACS also allows you to specify a number which is normally used to tell a command to execute many times. To repeat most commands, press META and then the number before you enter the command. Thus, the command META 5 ^F (M-5^F) will move the cursor forward by five characters. Try moving around in the text by using these commands. For extra practice, see how close you can come to the word ”small” in the first paragraph by giving an argument to the commands listed here.

Two other simple cursor commands that are useful to help us move around in the text are M-N next-paragraph which moves the cursor to the second paragraph, and M-P previous-paragraph which moves it back to the previous paragraph. The cursor may also be moved rapidly from one end of the line to the other. Move the cursor to the word ”few” in the second line. Press ^A beginning-of-line . Notice the cursor moves to the word ”events” at the beginning of the line. Pressing ^E end-of-line moves the cursor to the end of the line.

Finally, the cursor may be moved from any point in the file to the end or beginning of the file. Entering M-¿ end-of-file moves the cursor to the end of the buffer, M-< beginning-of-file to the first character of the file.

On the IBM-PC, the ATARI ST and many other machines, the cursor keys can also be used to move the cursor.

Practice moving the cursor in the text until you are comfortable with the commands we’ve explored in this chapter.

2.6 Saving your text

When you’ve finished practicing cursor movement, save your file. Your file currently resides in a BUFFER. The buffer is a temporary storage area for your text, and is lost when the computer is turned off. You can save the buffer to a file by entering ^X^S save-file . Notice that EMACS informs you that your file has no name and will not let you save it.

To save your buffer to a file with a different name than it’s current one (which is empty), press ^X^W write-file . EMACS will prompt you for the filename you wish to write. Enter the name fang.txt and press return. On a micro, the drive light will come on, and EMACS will inform you it is writing the file. When it finishes, it will inform you of the number of lines it has written to the disk.

Congratulations!! You’ve just saved your first EMACS file!

MicroEMACS Reference Manual  Basic Concepts

Chapter 2 Summary

In chapter 2, you learned how to enter text, how to use wrap mode, how to move the cursor, and to save a buffer. The following is a table of the commands covered in this chapter and their corresponding key bindings:

Key Binding  Keystroke Effect
abort-command  ^G  aborts current command
add-mode  ^XM  allows addition of EMACS mode such as WRAP
backward-character ^B  moves cursor left one character
beginning-of-file  M-¡  moves cursor to beginning of file
beginning-of-line  ^A  moves cursor to beginning of line
end-of-file  M-¿  moves cursor to end of file
end-of-line  ^E  moves cursor to end of line
forward-character ^F  moves cursor right one character
next-line  ^N  moves cursor to next line
next-paragraph  M-N  moves cursor to next paragraph
next-word  M-F  moves cursor forward one word
previous-line  ^P  moves cursor backward by one line
previous-paragraph M-P  moves cursor to previous paragraph
previous-word  M-B  moves cursor backward by one word
save-file  ^X^S  saves current buffer to a file
write-file  ^X^W  saves current buffer under a new name

Basic Editing–Simple Insertions and Deletions  MicroEMACS Reference Manual

Chapter 3

Basic Editing–Simple Insertions and Deletions

3.1 A Word About Windows, Buffers, Screens, and Modes

In the first chapter, you learned how to create and save a file in EMACS. Let’s do some more editing on this file. Call up emacs by typing in the following command.

emacs fang.txt

On icon oriented systems, double click on the uEMACS icon, usually a file dialog box of some sort will appear. Choose FANG.TXT from the appropriate folder.

Shortly after you invoke EMACS, the text should appear on the screen ready for you to edit. The text you are looking at currently resides in a buffer. A buffer is a temporary area of computer memory which is the primary unit internal to EMACS — this is the place where EMACS goes to work. The mode line at the bottom of the screen lists the buffer name, FANG.TXT and the name of the file with which this buffer is associated, FANG.TXT

The computer talks to you through the use of its screen. This screen usually has an area of 24 lines each of 80 characters across. You can use EMACS to subdivide the screen into several separate work areas, or windows, each of which can be ’looking into’ different files or sections of text. Using windows, you can work on several related texts at one time, copying and moving blocks of text between windows with ease. To keep track of what you are editing, each window is identified by a mode line on the last line of the window which lists the name of the buffer which it is looking into, the file from which the text was read, and how the text is being edited.

An EMACS mode tells EMACS how to deal with user input. As we have already seen, the mode ”WRAP” controls how EMACS deals with long lines (lines with over 79 characters) while the user is typing them in. The ”VIEW” mode, allows you to read a file without modifying it. Modes are associated with buffers and not with files; hence, a mode needs to be explicitly set or removed every time you edit a file. A new file read into a buffer with a previously specified mode will be edited under this mode. If you use specific modes frequently, EMACS allows you to set the modes which are used by all new buffers, called global modes.

3.2 Insertions

Your previously-saved text should look like this:

Fang Rock lighthouse, center of a series of mysterious and terrifying events at the turn of the century, is built on a rocky island a few miles of the Channel coast. So small is the island that wherever you stand its rocks are wet with sea spray. The lighthouse tower is in the center of the island. A steep flight of steps leads to the heavy door in its base. Winding stairs lead up to the crew room.

MicroEMACS Reference Manual  Basic Editing–Simple Insertions and Deletions

Let’s assume you want to add a sentence in the second paragraph after the word ”base.” Move the cursor until it is on the ”W” of ”Winding”. Now type the following:

This gives entry to the lower floor where the big steam generator throbs steadily away, providing power for the electric lantern.

If the line fails to wrap and you end up with a $ sign in the right margin, just enter M-Q fill-paragraph to reformat the paragraph. This new command attempts to fill out a paragraph. Long lines are divided up, and words are shuffled around to make the paragraph look nicer.

Notice that all visible EMACS characters are self-inserting — all you had to do was type the characters to insert and the existing text made space for it. With a few exceptions discussed later, all non-printing characters (such as control or escape sequences) are commands. To insert spaces, simply use the space bar. Now move to the first line of the file and type ^O open-line (Oh, not zero). You’ve just learned how to insert a blank line in your text.

3.3 Deletions

EMACS offers a number of deletion options. For example, move the cursor until it’s under the period at the end of the insertion you just did. Press the backspace key. Notice the ”n” on ”lantern” disappeared. The backspace implemented on EMACS is called a destructive backspace — it removes text immediately before the current cursor position from the buffer. Now type ^H delete-previous-character . Notice that the cursor moves back and obliterates the ”r” — either command will backspace the cursor.

Type in the two letters you erased to restore your text and move the cursor to the beginning of the buffer M-¿ beginning-of-file . Move the cursor down one line to the beginning of the first paragraph.

To delete the forward character, type ^D delete-next-character . The ”F” of ”Fang” disappears. Continue to type ^D until the whole word is erased EMACS also permits the deletion of larger elements of text. Move the cursor to the word ”center” in the first line of text. Pressing M-<backspace¿ delete-previous-word kills the word immediately before the cursor. M-^H has the same effect.

Notice that the commands are very similar to the control commands you used to delete individual letters. As a general rule in EMACS, control sequences affect small areas of text, META sequences larger areas. The word forward of the cursor position can therefore be deleted by typing M-D delete-next-word . Now let’s take out the remainder of the first line by typing ^K kill-to-end-of-line . You now have a blank line at the top of your screen. Typing ^K again or ^X^O deleteblank-lines deletes the blank line and flushes the second line to the top of the text. Now exit EMACS by typing ^X^C exit-emacs . Notice EMACS reminds you that you have not saved your buffer. Ignore the warning and exit. This way you can exit EMACS without saving any of the changes you just made.

Basic Editing–Simple Insertions and Deletions  MicroEMACS Reference Manual

Chapter 3 Summary

In Chapter 3, you learned about the basic ”building blocks” of an EMACS text file — buffers, windows, and files.

Key binding  Keystroke Effect
delete-previous-character ^H  deletes character immediately before the current cursor position
delete-next-character  ^D  deletes character immediately after current cursor position
delete-previous-word  M-^H  deletes word immediately before current cursor position
delete-next-word  M-D  deletes word immediately after current cursor position
kill-to-end-of-line  ^K  deletes from current cursor position to end of line
insert-space  ^C  inserts a space to right of cursor
open-line  ^O  inserts blank line
delete-blank-lines  ^X^O  removes blank line
exit-emacs  ^X^C  exits emacs

10

Using Regions  MicroEMACS Reference Manual

Chapter 4

Using Regions

4.1 Defining and Deleting a Region

At this point its time to familiarize ourselves with two more EMACS terms — the point and the mark. The point is located directly behind the current cursor position. The mark (as we shall see shortly) is user defined. These two elements together are called the current region and limit the region of text on which EMACS performs many of its editing functions.

Let’s begin by entering some new text. Don’t forget to add wrap mode if its not set on this buffer. Start EMACS and open a file called PUBLISH.TXT. Type in the following text:

One of the largest growth areas in personal computing is electronic publishing. There are packages available for practically every machine from elegantly simple programs for the humble Commodore 64 to sophisticated professional packages for PC and Macintosh computers. Electronic publishing is as revolutionary in its way as the Gutenburg press. Whereas the printing press allowed the mass production and distribution of the written word, electronic publishing puts the means of production in the hands of nearly every individual. From the class magazine to the corporate report, electronic publishing is changing the way we produce and disseminate information. Personal publishing greatly increases the utility of practically every computer. Thousands of people who joined the computer revolution of this decade only to hide their machines unused in closets have discovered a new use for them as dedicated publishing workstations.

Now let’s do some editing. The last paragraph seems a little out of place. To see what the document looks like without it we can cut it from the text by moving the cursor to the beginning of the paragraph. Enter M-<space¿ set-mark . EMACS will respond with [Mark set]. Now move the cursor to the end of the paragraph. You have just defined a region of text. To remove this text from the screen, type ^W kill-region . The paragraph disappears from the screen.

On further consideration, however, perhaps the paragraph we cut wasn’t so bad after all. The problem may have been one of placement. If we could tack it on to the end of the first paragraph it might work quite well to support and strengthen the argument. Move the cursor to the end of the first paragraph and enter ^Y yank . Your text should now look like this:

One of the largest growth areas in personal computing is electronic publishing. There are packages available for practically every machine from elegantly simple programs for the humble Commodore 64 to sophisticated professional packages for PC and Macintosh computers. Personal publishing greatly increases the utility of practically every computer. Thousands of people who joined the computer revolution of this decade only to hide their machines unused in closets have discovered a new use for them as dedicated publishing workstations. Electronic publishing is as revolutionary in its way as the Gutenburg press. Whereas the printing press allowed the mass production and distribution of the written word, electronic publishing puts the means of production in the hands of nearly every individual. From the class magazine to the corporate report, electronic publishing is changing the way we produce and disseminate information.

MicroEMACS Reference Manual  Using Regions

4.2 Yanking a Region

The text you cut initially didn’t simply just disappear, it was cut into a buffer that retains the ”killed” text appropriately called the kill buffer. ^Y ”yanks” the text back from this buffer into the current buffer. If you have a long line (indicated, remember, by the $ sign), simply hit M-Q to reformat the paragraph.

There are other uses to which the kill buffer can be put. Using the method we’ve already learned, define the last paragraph as a region. Now type M-W copy-region . Nothing seems to have happened; the cursor stays blinking at the point. But things have changed, even though you may not be able to see any alteration.

To see what has happened to the contents of the kill buffer, move the cursor down a couple of lines and ”yank” the contents of the kill buffer back with ^Y. Notice the last paragraph is now repeated. The region you defined is ”tacked on” to the end of your file because M-W copies a region to the kill buffer while leaving the original text in your working buffer. Some caution is needed however, because the contents of the kill buffer are updated when you delete any regions, lines or words. If you are moving large quantities of text, complete the operation before you do any more deletions or you could find that the text you want to move has been replaced by the most recent deletion. Remember — a buffer is a temporary area of computer memory that is lost when the machine is powered down or switched off. In order to make your changes permanent, they must be saved to a file before you leave EMACS. Let’s delete the section of text we just added and save the file to disk.

Chapter 4 Summary

In Chapter 4, you learned how to achieve longer insertions and deletions. The EMACS terms point and mark were introduced and you learned how to manipulate text with the kill buffer.

Key Binding  Keystroke  Effect
set-mark  M-<space¿ Marks the beginning of a region
kill-region  ^W  Deletes region between point and mark and places it in KILL buffer
copy-region  M-W  Copies text between point and mark into KILL buffer
yank  ^Y  Inserts a copy of the KILL buffer into current buffer at point
yank-pop  M-W  Acts as @t(yank) the first time, but if repeated without moving the cursor the next newest killed text replaces the text just yanked (up to 16 kills saved)
cycle-ring  ^X-Y  Moves the current position in the ring
delete-kill-ring M-^Y  Dumps all the killed text in the ring

Search and Replace  MicroEMACS Reference Manual

Chapter 5

Search and Replace

5.1 Forward Search

Load EMACS and bring in the file you just saved. Your file should look like the one below.

One of the largest growth areas in personal computing is electronic publishing. There are packages available for practically every machine from elegantly simple programs for the humble Commodore 64 to sophisticated professional packages for PC and Macintosh computers. Personal publishing greatly increases the utility of practically every computer. Thousands of people who joined the computer revolution of this decade only to hide their machines unused in closets have discovered a new use for them as dedicated publishing workstations. Electronic publishing is as revolutionary in its way as the Gutenburg press. Whereas the printing press allowed the mass production and distribution of the written word, electronic publishing puts the means of production in the hands of nearly every individual. From the class magazine to the corporate report, electronic publishing is changing the way we produce and disseminate information.

Let’s use EMACS to search for the word ”revolutionary” in the second paragraph. Because EMACS searches from the current cursor position toward the end of buffers, and we intend to search forward, move the cursor to the beginning of the text. Enter ^S search-forward . Note that the command line now reads

     Search [] <META>:

EMACS is prompting you to enter the search string — the text you want to find. Enter the word revolutionary and hit the META key. The cursor moves to the end of the word ”revolutionary.”

Notice that you must enter the <META> key to start the search. If you simply press <NL> the command line responds with ‘<NL>’. Although this may seem infuriating to users who are used to pressing the return key to execute any command, EMACS’ use of <META> to begin searches allows it to pinpoint text with great accuracy. After every line wrap or carriage return, EMACS ”sees” a new line character (<NL>). If you need to search for a word at the end of a line, you can specify this word uniquely in EMACS.

In our sample text for example, the word ”and” occurs a number of times, but only once at the end of a line. To search for this particular occurrence of the word, move the cursor to the beginning of the buffer and type ^S. Notice that EMACS stores the last specified search string as the default string. If you press <META¿ now, EMACS will search for the default string, in this case, ”revolutionary”.

To change this string so we can search for our specified ”and” simply enter the word and followed by <NL¿. The command line now shows:

     Search [and<NL>]<META>:

Press <META¿ and the cursor moves to ‘and’ at the end of the second last line.

MicroEMACS Reference Manual  Search and Replace

5.2 Exact Searches

If the mode EXACT is active in the current buffer, EMACS searches on a case sensitive basis. Thus, for example you could search for Publishing as distinct from publishing.

5.3 Backward Search

Backward searching is very similar to forward searching except that it is implemented in the reverse direction. To implement a reverse search, type ^R search-reverse . Because EMACS makes no distinction between forward and backward stored search strings, the last search item you entered appears as the default string. Try searching back for any word that lies between the cursor and the beginning of the buffer. Notice that when the item is found, the point moves to the beginning of the found string (i.e., the cursor appears under the first letter of the search item).

Practice searching for other words in your text.

5.4 Searching and Replacing

Searching and replacing is a powerful and quick way of making changes to your text. Our sample text is about electronic publishing, but the correct term is ”desktop” publishing. To make the necessary changes we need to replace all occurrences of the word ”electronic” with ”desktop.” First, move the cursor to the top of the current buffer with the M-< command. Then type M-R replace-string . The command line responds:

     Replace []<META>:

where the square brackets enclose the default string. Type the word electronic and hit <META¿. The command line responds:

     with []<META>

type desktop<META¿. EMACS replaces all instances of the original word with your revision. Of course, you will have to capitalize the first letter of ”desktop” where it occurs at the beginning of a sentence.

You have just completed an unconditional replace. In this operation, EMACS replaces every instance of the found string with the replacement string.

5.5 Query-Replace

You may also replace text on a case by case basis. The M-^R query-replace-string command causes EMACS to pause at each instance of the found string.

For example, assume we want to replace some instances of the word ”desktop” with the word ”personal.” Go back to the beginning of the current buffer and enter the M-^R query-replace-string command. The procedure is very similar to that which you followed in the unconditional search/replace option. When the search begins however, you will notice that EMACS pauses at each instance of ”publishing” and asks whether you wish to replace it with the replacement string. You have a number of options available for response:

Search and Replace  MicroEMACS Reference Manual
Response  Effect
Y(es)  Make the current replacement and skip to the next occurrence of the search string
N(o)  Do not make this replacement but continue
!  Do the rest of the replacements with no more queries
U(ndo)  Undo just the last replacement and query for it again (This can only go back ONE time)
^G  Abort the replacement command (This action does not undo previously-authorized replacements
.  Same effect as ^G, but cursor returns to the point at which the replacement command was given
?  This lists help for the query replacement command

Practice searching and searching and replacing until you feel comfortable with the commands and their effects.

Chapter 5 Summary

In chapter 5, you learned how to search for specified strings of text in EMACS. The chapter also dealt with searching for and replacing elements within a buffer.

Key Binding  Keystroke Effect

search-forward ^S Searches from point to end of buffer. Point is moved from current location to the end

of the found string

search-backward ^R Searches from point to beginning of buffer. Point is moved from current location to

beginning of found string

replace M-R Replace ALL occurrences of search string with specified (null) string from point to

the end of the current buffer

query-replace-string M-^R  As above, but pause at each found string and query for action

16

Windows  MicroEMACS Reference Manual

Chapter 6

Windows

6.1 Creating Windows

We have already met windows in an earlier chapter. In this chapter, we will explore one of EMACS’ more powerful features — text manipulation through multiple windowing. Windows offer you a powerful and easy way to edit text. By manipulating a number of windows and buffers on the screen simultaneously, you can perform complete edits and revisions on the computer screen while having your draft text or original data available for reference in another window.

You will recall that windows are areas of buffer text that you can see on the screen. Because EMACS can support several screen windows simultaneously you can use them to look into different places in the same buffer. You can also use them to look at text in different buffers. In effect, you can edit several files at the same time.

Let’s invoke EMACS and pull back our file on desktop publishing by typing

emacs publish.txt

When the text appears, type the ^X2 split-current-window command. The window splits into two windows. The window where the cursor resides is called the current window — in this case the bottom window. Notice that each window has a text area and a mode line. The command line is however, common to all windows on the screen.

The two windows on your screen are virtually mirror images of each other because the new window is opened into the same buffer as the one you are in when you issue the open-window command . All commands issued to EMACS are executed on the current buffer in the current window.

To move the cursor to the upper window (i.e., to make that window the current window, type ^XP previous-window . Notice the cursor moves to the upper or previous window. Entering ^XO next-window moves to the next window. Practice moving between windows. You will notice that you can also move into the Function Key menu by entering these commands.

Now move to the upper window. Let’s open a new file. On the EMACS disk is a tutorial file. Let’s call it into the upper window by typing:

     ^X^F

and press return.

Enter the filename emacs.tut.

In a short time, the tutorial file will appear in the window. We now have two windows on the screen, each looking into different buffers. We have just used the ^X^F find-file command to find a file and bring it into our current window.

You can scroll any window up and down with the cursor keys, or with the commands we’ve learned so far. However, because the area of visible text in each window is relatively small, you can scroll the current window a line at a time.

MicroEMACS Reference Manual  Windows

Type ^X^N move-window-down The current window scrolls down by one line — the top line of text scrolls out of view, and the bottom line moves towards the top of the screen. You can imagine, if you like, the whole window slowly moving down to the end of the buffer in increments of one line. The command ^X^P move-window-up scrolls the window in the opposite direction.

As we have seen, EMACS editing commands are executed in the current window, but the program does support a useful feature that allows you to scroll the next window. M-^Z scroll-next-up scrolls the next window up, M-^V scroll-next-down scrolls it downward. From the tutorial window, practice scrolling the window with the desktop publishing text in it up and down.

When you’re finished, exit EMACS without saving any changes in your files.

Experiment with splitting the windows on your screen. Open windows into different buffers and experiment with any other files you may have. Try editing the text in each window, but don’t forget to save any changes you want to keep — you still have to save each buffer separately.

6.2 Deleting Windows

Windows allow you to perform complex editing tasks with ease. However, they become an inconvenience when your screen is cluttered with open windows you have finished using. The simplest solution is to delete unneeded windows. The command ^X0 delete-window will delete the window you are currently working in and move you to the next window.

If you have a number of windows open, you can delete all but the current window by entering ^X1 delete-other-windows.

6.3 Resizing Windows

During complex editing tasks, you will probably find it convenient to have a number of windows on the screen simultaneously. However this situation may present inconveniences because the more windows you have on the screen the smaller they are; in some cases, a window may show only a couple of lines of text. To increase the flexibility and utility of the window environment, EMACS allows you to resize the window you are working in (called, as you will recall, the current window) to a convenient size for easier editing, and then shrink it when you no longer need it to be so large.

Let’s try an example. Load in any EMACS text file and split the current window into two. Now type ^X^(Shift-6), growwindow. Your current window should be the lower one on the screen. Notice that it increases in size upwards by one line. If you are in the upper window, it increases in size in a downward direction. The command ^X^Z, shrink-window correspondingly decreases window size by one line at a time.

EMACS also allows you to resize a window more precisely by entering a numeric argument specifying the size of the window in lines. To resize the window this way, press the META key and enter a numeric argument (remember to keep it smaller than the number of lines on your screen display) then press ^XW resize-window. The current window will be enlarged or shrunk to the number of lines specified in the numeric argument. For example entering:

M-8 ^XW

will resize the current window to 8 lines.

6.4 Repositioning within a Window

The cursor may be centered within a window by entering M-! or M-^L redraw-display . This command is especially useful in allowing you to quickly locate the cursor if you are moving frequently from window to window. You can also use this command to move the line containing the cursor to any position within the current window. This is done by using a numeric argument before the command. Type M-<n¿ M-^L where <n> is the number of the line within the window that you wish the current line to be displayed.

The ^L clear-and-redraw command is useful for ”cleaning up” a ”messy” screen that can result of using EMACS on a mainframe system and being interrupted by a system message.

Windows  MicroEMACS Reference Manual

Chapter 6 Summary

In Chapter 6, you learned how to manipulate windows and the editing flexibility they offer.

Key Binding  Keystroke Effect
open-window  ^X2  Splits current window into two windows if space available
close-windows  ^X1  Closes all windows except current window
next-window  ^XO[oh] Moves point into next (i.e. downward) window
previous-window ^XP Moves point to previous (i.e. upward) window
move-window-down ^X^N  Scrolls current window down one line
move-window-up  ^X^P  Scrolls current window up one line
redraw-display  M ! or  Window is moved so line with point
M ^L  (with cursor) is at center of window

grow-window M-X ^Current window is enlarged by one line and nearest window is shrunk

by one line

shrink-window ^X^Z Current window is shrunk by one line and nearest window is enlarged

by one line

clear-and-redraw ^L Screen is blanked and redrawn. Keeps screen updates in sync with your

commands

scroll-next-up  M-^Z  Scrolls next window up by one line
scroll-next-down  M-^V  Scrolls next window down by one line
delete-window  ^X0  Deletes current window
delete-other-windows ^X1  Deletes all but current window
resize-window  ^X^W  Resizes window to a given numeric argument

20

Using a Mouse  MicroEMACS Reference Manual

Chapter 7

Using a Mouse

On computers equipped with a mouse, the mouse can usually be used to make editing easier. If your computer has a mouse, let’s try using it. Start MicroEMACS by typing:

emacs publish.txt

This brings EMACS up and allows it to edit the file from the last chapter. If the function key window is visible on the screen, press the F5 key to cause it to disappear. Now use the ^X2 split-current-window command to split the screen into two windows. Next use the ^X^F find-file command to read in the fang.txt file. Now your screen should have two windows looking into two different files.

Grab the mouse and move it around. On the screen an arrow, or block of color appears. This is called the mouse cursor and can be positioned on any character on the screen. On some computers, positioning the mouse cursor in the extreme upper right or left corner may bring down menus which allow you to access that computers utilities, sometimes called Desk Accessories .

7.1 Moving around with the mouse

Using the mouse button (or the left button if the mouse has more than one), position the mouse over some character in the current window. Click the mouse button once. The point will move to where the mouse cursor is. If you place the mouse cursor past the end of a line, the point will move to the end of that line.

Move the mouse cursor into the other window and click on one of the characters there. MicroEMACS will automatically make this window the current window (notice that the mode line changes) and position the point to the mouse cursor. This makes it very easy to use the mouse to switch to a different window quickly.

7.2 Dragging around

Besides just using the mouse to move around on the screen, you can use the same button to move text. Move the mouse cursor to a character in one of the windows, and click down... but don’t let the button up yet! The point will move to where the mouse cursor is. Now move the mouse cursor up or down on the screen, and release the button. The point will again move to where the mouse cursor is, but this time it will bring the text under it along for the ride. This is called dragging, and is how you can make the text appear just where you want it to. If you try to drag text out of the current window, EMACS will ignore your attempt and leave the point where you first clicked down.

Now, click down on a word in one of the windows, and drag it directly to the left. Release the button and watch as the entire window slides, or scrolls to the left. The missing text has not been deleted, it is simply not visible, off the left hand side of the screen. Notice the mode line has changed and now looks like:

MicroEMACS Reference Manual  Using a Mouse

==== MicroEMACS 3.12 [<12] () == fang.txt == File:  fang.txt ==============

The number insided the brackets [] shows that the screen is now scrolled 12 characters from the left margin.

Now grab the same text again, and drag it to the right, pulling the rest of the text back into the current window. The [<] field will disappear, meaning that the window is no longer scrolled to the left. This feature is very useful for looking at wide charts and tables. Remember, MicroEMACS will only scroll the text in the current window sideways if you drag it straight to the side, otherwise it will drag the text vertically.

Now, place the mouse cursor over a character on the upper mode line, click down, move the mouse cursor up or down a few lines and let go of the button. The mode line moves to where you dragged it, changing the size of the windows above and below it. If you try to make a window with less than one line, EMACS will not let you. Dragging the mode lines can make it very fast and easy for you to rearrange the windows as you would like.

If you have a number of different windows visible on the screen, positioning the mouse over the mode line of one window and clicking the right mouse button will cause that window to be deleted.

7.3 Cut and Paste

If your mouse has two buttons, then you can use the right button to do some other things as well. Earlier, we learned how to define a region by using the M-<space¿ set-mark command. Now, position the mouse over at the beginning of a region you would like to copy. Next click and hold down the right mouse button. Notice that the point jumps to the mouse cursor and EMACS reports [Mark Set]. Holding the button down move the mouse to the end of the text you wish to copy and release the mouse button. Emacs reports [Region Copied] to let you know it has copied the region into the KILL buffer. This has done the same job as the M-W copy-region command.

If you now click the right mouse button, without moving the mouse, the region you defined disappear, being cut from the current buffer. This works just like the ^W kill-region command.

If you move the mouse away from where you cut the text, and click the right mouse button down and up without moving the mouse, the text in the KILL buffer gets inserted, or pasted into the current buffer at the point.

7.4 Screens

MicroEMACS can use more than one screen at once. Each screen is a collection of windows along with a mode line. These screens usually fill the terminal or computer screen on text based systems, but can also be held in different windows on graphically based systems like MicroSoft Windows, OS/2, the Macintosh Finder and X-Windows. Don’t be confused by the two different uses of the term ”window”. Inside EMACS style editors, a window lets you view part of a buffer. Under graphical operating systems, a window holds a ”virtual terminal”, allowing you to manipulate more than one job, editing session or program at once. Within MicroEMACS, these operating system windows are called screens. All these screens are displayed on your current desktop.

7.5 Resizing a Screen

You can change the size of a screen. Move the mouse to the last position of the command line. Press the left mouse button down. Holding it, move the mouse to the place you want the new lower right corner. Release the mouse. The desktop redraws, with your newly resized screen. MicroEMACS will ignore size changes that can not be done, like attempting to pull the lower left corner above the upper right corner of the current screen.

7.6 Moving a Screen

To change where on the desktop a screen is placed, move the mouse to the upper right corner of the screen, press the left mouse button down, move the mouse and release it where you want the screen displayed. Again, MicroEMACS will ignore placements that can not be done.

7.7 Creating a Screen

Creating a new screen is just like moving a screen, but using the right button. Move to the upper right of an existing screen, press the right mouse button down, and move the mouse, releasing the button where the new screen should appear. A new screen will have a single window, containing the contents of the current window in the copied screen, and will have that window’s colors. The new screen will have the copied screen’s size.

7.8 Switching to a Screen

This is simple. Any mouse command can be done in any screen by placing the mouse on a visible part of the screen and clicking. The last screen the mouse is used on comes to front and is the current screen. Also, the A-C cycle-screens command brings the rearmost screen to front.

7.9 Deleting a Screen

Place the mouse on the command line of the screen you want to delete. Click the right mouse button, the screen will disappear. If you delete the only remaining screen on the desktop, MicroEMACS will exit.

Chapter 7 Summary

In Chapter 7, you learned how to use the mouse to move the point, switch windows, drag text, and resize windows. You also learned how to use the right mouse button in order to copy and delete regions and yank them back at other places. And lastly, you learned how to control multiple screens with the mouse.

Action  Mouse Directions
Move Cursor  position mouse cursor over desired location click down and up with left button

Drag Text position mouse cursor over desired text, click left button down, move to new screen location

for text, release mouse button

Resize Windows position mouse cursor over mode line to move click left button down, move to new location

for mode line, release mouse button

Delete Window position mouse cursor over mode line of window to delete, click right mouse button
Activate Screen move mouse over existing screen, click left button down and up

Resize Screen position mouse cursor over last character on message line, click left button down, move to

new lower right corner of screen, release mouse button

Copy Region position mouse at beginning of region, click right button down, move to end of region, release

mouse button

Cut Region position mouse at beginning of region, click right button down, move to end of region, release

mouse button, click right button down and up

Paste Region  position mouse at place to paste, click right button down and up

Create Screen position mouse at upper left corner of existing screen, click right button down, move to posi-

tion of new screen, release mouse button

Resize Screen position mouse at lower right corner of screen, click left button down, move to new lower left

corner, release mouse button

Move Screen position mouse at upper right corner of screen, click left button down, move to new screen

position, release mouse button

Delete Screen  position to command line of existing screen, click right button down, release mouse button

24

Buffers  MicroEMACS Reference Manual

Chapter 8

Buffers

We have already learned a number of things about buffers. As you will recall, they are the major internal entities in EMACS — the place where editing commands are executed. They are characterized by their names, their modes, and by the file with which they are associated. Each buffer also ”remembers” its mark and point. This convenient feature allows you to go to other buffers and return to the original location in the ”current” buffer.

Advanced users of EMACS frequently have a number of buffers in the computer’s memory simultaneously. In the last chapter, for example, you opened at least two buffers — one into the text you were editing, and the other into the EMACS on-line tutorial. If you deal with complex text files — say, sectioned chapters of a book, you may have five or six buffers in the computer’s memory. You could select different buffers by simply calling up the file with ^X^F find-file , and let EMACS open or reopen the buffer. However, EMACS offers fast and sophisticated buffering techniques that you will find easy to master and much more convenient to use.

Let’s begin by opening three buffers. You can open any three you choose, for example call the following files into memory: fang.txt, publish.txt, and emacs.tut in the order listed here. When you’ve finished this process, you’ll be looking at a screen showing the EMACS tutorial. Let’s assume that you want to move to the fang.txt buffer. Enter:

^XX next-buffer

This command moves you to the next buffer. Because EMACS cycles through the buffer list, which is alphabetized, you will now be in the fang.txt buffer. Using ^XX again places you in the publish.txt buffer. If you are on a machine that supports function keys, using ^XX again places you in the Function Keys buffer. Using ^XX one last time cycles you back to the beginning of the list.

If you have a large number of buffers to deal with, this cycling process may be slow and inconvenient. The command ^XB select-buffer allows you to specify the buffer you wish to be switched to. When the command is entered, EMACS prompts, Use buffer:. Simply enter the buffer name (NOT the file name), and that buffer will then become the current buffer. If you type in part of the file name and press the space bar, EMACS will attempt to complete the name from the list of current buffers. If it succeeds, it will print the rest of the name and you can hit <NL> to switch to that buffer. If EMACS beeps the bell, there is no such buffer, and you may continue editing the name on the command line.

Multiple buffer manipulation and editing is a complex activity, and you will probably find it very inconvenient to re-save each buffer as you modify it. The command ^X^B list-buffers creates a new window that gives details about all the buffers currently known to EMACS. Buffers that have been modified are identified by the ”buffer changed” indicator (an asterisk in the second column). You can thus quickly and easily identify buffers that need to be saved to files before you exit EMACS. The buffer window also provides other information — buffer specific modes, buffer size, and buffer name are also listed. To close this window, simply type the close-windows command, ^X1.

To delete any buffer, type ^XK delete-buffer . EMACS prompts you Kill buffer:. Enter the buffer name you want to delete. As this is destructive command, EMACS will ask for confirmation if the buffer was changed and not saved. Answer Y(es) or N(o). As usual ^G cancels the command.

MicroEMACS Reference Manual  Buffers

Chapter 8 Summary

In Chapter 8, you learned how to manipulate buffers.

Key Binding Keystroke Effect
next-buffer ^X^X Switch to the next buffer in the buffer list
select-buffer ^XB  Switch to a particular buffer
list-buffers ^X^B List all buffers
delete-buffer ^XK  Delete a particular buffer if it is off-screen

Modes  MicroEMACS Reference Manual

Chapter 9

Modes

EMACS allows you to change the way it works in order to customized it to the style of editing you are using. It does this by providing a number of different modes . These modes can effect either a single buffer, or any new buffer that is created. To add a mode to the current buffer, type ^XM add-mode . EMACS will then prompt you for the name of a mode to add. When you type in a legal mode name, and type a <NL>, EMACS will add the mode name to the list of current mode names in the mode line of the current buffer.

To remove an existing mode, typing the ^X^M delete-mode will cause EMACS to prompt you for the name of a mode to delete from the current buffer. This will remove that mode from the mode list on the current mode line.

Global modes are the modes which are inherited by any new buffers which are created. For example, if you wish to always do string searching with character case being significant, you would want global mode EXACT to be set so that any new files read in inherent the EXACT mode. Global modes are set with the M-M add-global-mode command, and unset with the M-^M delete-global-mode command. Also, the current global modes are displayed in the first line of a ^X^B list-buffers command.

On machines which are capable of displaying colors, the mode commands can also set the background and foreground character colors. Using add-mode or delete-mode with a lowercase color will set the background color in the current window. An uppercase color will set the foreground color in the current window. Colors that EMACS knows about are: white, cyan, magenta, yellow, blue, red, green, and black. If the computer you are running on does not have eight colors, EMACS will attempt to make some intelligent guess at what color to use when you ask for one which is not there.

9.1 ASAVE mode

Automatic Save mode tells EMACS to automatically write out the current buffer to its associated file on a regular basis. Normally this will be every 256 characters typed into the file. The environment variable $acount counts down to the next auto-save, and $asave is the value used to reset $acount after a save occurs.

9.2 CMODE mode

CMODE is useful to C programmers. When CMODE is active, EMACS will try to assist the user in a number of ways. This mode is set automatically with files that have a .c or .h extension.

The <NL> key will normally attempt to return the user to the next line at the same level of indentation as the last non blank line, unless the current line ends with a open brace ({) in which case the new line will be further indented by one tab position.

A close brace (}) will search for the corresponding open brace and line up with it.

A pound sign (#) with only leading white space will delete all the white space preceding itself. This will always bring preprocessor directives flush to the left margin.

MicroEMACS Reference Manual  Modes

Whenever any close fence is typed, IE )]>}, if the matching open fence is on screen in the current window, the cursor will briefly flash to it, and then back. This makes balancing expressions, and matching blocks much easier.

9.3 CRYPT mode

When a buffer is in CRYPT mode, it is encrypted whenever it is written to a file, and decrypted when it is read from the file. The encryption key can be specified on the command line with the -k switch, or with the M-E set-encryption-key command. If you attempt to read or write a buffer in crypt mode and now key has not been set, EMACS will execute set-encryption-key automatically, prompting you for the needed key. Whenever EMACS prompts you for a key, it will not echo the key to your screen as you type it (IE make SURE you get it right when you set it originally).

The encryption algorithm used changes all characters into normal printing characters, thus the resulting file is suitable for sending via electronic mail. All version of MicroEMACS should be able decrypt the resulting file regardless of what machine encrypted it. Also available with EMACS is the stand alone program, MicroCRYPT, which can en/decrypt the files produced by CRYPT mode in EMACS.

9.4 EXACT mode

All string searches and replacements will take upper/lower case into account. Normally the case of a string during a search or replace is not taken into account.

9.5 MAGIC mode

In the MAGIC mode certain characters gain special meanings when used in a search pattern. Collectively they are know as regular expressions, and a limited number of them are supported in MicroEmacs. They grant greater flexibility when using the search command. They have no affect on the incremental search command.

The symbols that have special meaning in MAGIC mode are ^, $, ., &, ?, *, +, [ (and ], used with it), and \.

The characters ^and $ fix the search pattern to the beginning and end of line, respectively. The ^character must appear at the beginning of the search string, and the $ must appear at the end, otherwise they lose their meaning and are treated just like any other character. For example, in MAGIC mode, searching for the pattern t$ would put the cursor at the end of any line that ended with the letter t. Note that this is different than searching for t<NL>, that is, t followed by a newline character. The character $ (and ^, for that matter) matches a position, not a character, so the cursor remains at the end of the line. But a newline is a character that must be matched like any other character, which means that the cursor is placed just after it - on the beginning of the next line.

The character . has a very simple meaning — it matches any single character, except the newline. Thus a search for bad.er could match badger, badder (slang), or up to the r of bad error.

The character [ indicates the beginning of a character class. It is similar to the ”any” character ., but you get to choose which characters you want to match. The character class is ended with the character ]. So, while a search for ba.e will match bane, bade, bale, bate, et cetera, you can limit it to matching babe and bake by searching for ba[bk]e. Only one of the characters inside the [ and ] will match a character. If in fact you want to match any character except those in the character class, you can put a ^as the first character. It must be the first character of the class, or else it has no special meaning. So, a search for [^aeiou] will match any character except a vowel, but a search for [aeiou^] will match any vowel or a ^.

If you have many characters in order, that you want to put in the character class, you may use a dash (-) as a range character. So, [a-z] will match any letter (or any lower case letter if EXACT mode is on), and [0-9a-f] will match any digit or any letter a through f, which happen to be the characters for hexadecimal numbers. If the dash is at the beginning or end of a character class, it is taken to be just a dash.

The ? character indicates that the preceding character is optional. The character may or may not appear in the matched string. For example, a search for bea?st would match both beast and best. If there is no preceding charcter for ? to modify, it is treated as a normal question mark character.

The * character is known as closure, and means that zero or more of the preceding character will match. If there is no preceding character, * has no special meaning and is treated as a normal asterisk. The closure symbol will also have no special meaning if it is preceded by the beginning of line symbol ^, since it represents a position, not a character.

The notion of zero or more characters is important. If, for example, your cursor was on the line

This line is missing two vowels.

and a search was made for a*, the cursor would not move, because it is guaranteed to match no letter a , which satisfies the search conditions. If you wanted to search for one or more of the letter a, you could search for aa*, which would match the letter a, then zero or more of them. A better way, however, is to use the + character.

The + character behaves in every respect like the * character, with the exception that its minimum match range is one, not zero. Thus the pattern a+ is identical to aa*.

Under older versions of MicroEMACS, the closure symbols would not operate on newlines. The current versions no longer have this restriction.

The \ is the escape character. With the exception of groups, which are explained below, the \ is used at those times when you want to be in MAGIC mode, but also want a regular expression character to be just a character. It turns off the special meaning of the character. So a search for it\. will search for a line with it., and not it followed by any other character. Or, a search for TEST\*+ would match the word TEST followed by one or more asterisks. The escape character will also let you put ^, -, or ] inside a character class with no special side effects.

The character pair \( represent the start of a group in a search string. A group is ended by the character pair \). All characters matched within the \( and \) are part of a numbered group, and may be referenced with the &GROUP function, or with a \ followed by the group number in the replacement string of replace-string or the query-replace-string commands. For example, a search for INDEX\([0-9]+\), to be replaced by getind(\1) would change

indptr := INDEX42

to

indptr := getind(42)

There may be up to nine groups. Groups may be nested.

The character & (ampersand) is a replacement character, and represents all the characters which were matched by the search string. When used in the M-R replace-string or the M-^R query-replace-string commands, the & will be substituted for the search string.

9.6 OVER mode

OVER mode stands for overwrite mode. When in this mode, when characters are typed, instead of simply inserting them into the file, EMACS will attempt to overwrite an existing character past the point. This is very useful for adjusting tables and diagrams.

9.7 WRAP mode

Wrap mode is used when typing in continuous text. Whenever the cursor is past the currently set fill column (72 by default) and the user types a space or a <NL>, the last word of the line is brought down to the beginning of the next line. Using this, one just types a continuous stream of words and EMACS automatically inserts <NL>s at appropriate places.

NOTE to programmers:

The EMACS variable $wraphook contains the name of the function which executes when EMACS detects it is time to wrap. This is set to the function wrap-word by default, but can be changed to activate different functions and macroes at wrap time.

9.8 VIEW mode

VIEW mode disables all commands which can change the current buffer. EMACS will display an error message and ring the bell every time you attempt to change a buffer in VIEW mode.

Chapter 9 Summary

In Chapter 9, you learned about modes and their effects.

Key Binding  Keystroke Effect
add-mode  ^XM  Add a mode to the current buffer
delete-mode  ^X^M  Delete a mode from the current buffer
add-global-mode  M-M  Add a global mode to the current buffer
delete-global-mode M-^M  Delete a global mode from the current buffer

Files  MicroEMACS Reference Manual

Chapter 10

Files

A file is simply a collection of related data. In EMACS we are dealing with text files — named collections of text residing on a disk (or some other storage medium). You will recall that the major entities EMACS deals with are buffers. Disk-based versions of files are only active in EMACS when you are reading into or writing out of buffers. As we have already seen, buffers and physical files are linked by associated file names. For example, the buffer ”ch7.txt” which is associated with the physical disk file ch7.txt. You will notice that the file is usually specified by the drive name or (in the case of a hard drive) a path. Thus you can specify full file names in EMACS,

e.g. disk:\directories\filename.extension

If you do not specify a disk and directories, the default disk and the current directory is used.

IMPORTANT — If you do not explicitly save your buffer to a file, all your edits will be lost when you leave EMACS (although EMACS will prompt you when you are about to lose edits by exiting). In addition, EMACS does not protect your disk-based files from overwriting when it saves files. Thus when you instruct EMACS to save a file to disk, it will create a file if the specified file doesn’t exist, or it will overwrite the previously saved version of the file thus replacing it. Your old version is gone forever.

If you are at all unsure about your edits, or if (for any reason) you wish to keep previous versions of a file, you can change the name of the associated file with the command ^XN change-file-name. When this file is saved to disk, EMACS will create a new physical file under the new name. The earlier disk file will be preserved.

For example, let’s load the file fang.txt into EMACS. Now, type ^XN. The EMACS command line prompts Name:. Enter a new name for the file — say new.txt and press <NL>. The file will be saved under the new filename, and your disk directory will show both fang.txt and new.txt.

An alternative method is to write the file directly to disk under a new filename. Let’s pull our publish.txt file into EMACS. To write this file under another filename, type ^X^W write-file. EMACS will prompt you write file:. Enter an alternate filename — desktop.txt. Your file will be saved as the physical file desktop.txt.

Note that in the examples above, although you have changed the names of the related files, the buffer names remain the same. However, when you pull the physical file back into EMACS, you will find that the buffer name now relates to the filename.

For example — You are working with a buffer fang.txt with the related file fang.txt. You change the name of the file to new.txt. EMACS now shows you working with the buffer fang.txt and the related file new.txt. Now pull the file new.txt into EMACS. Notice that the buffer name has now changed to ‘new.txt.

If for any reason a conflict of buffer names occurs, (if you have files of the same name on different drives for example) EMACS will prompt you use buffer:. Enter an alternative buffer name if you need to.

For a list of file related commands (including some we‘ve already seen), see the summary page.

MicroEMACS Reference Manual  Files

Chapter 10 Summary

In Chapter 10, you learned some of the more advanced concepts of file naming and manipulation. The relationship between files and buffers was discussed in some detail.

Key Binding  Keystroke Effect
save-file  ^X^S  Saves contents of current buffer with associated filename on default disk/ di-

rectory (if not specified)

write-file ^X^W Current buffer contents will be saved under specified name

change-file-name ^XN The associated filename is changed (or associated if not previously specified)

as specified

find-file ^X^F Reads specified file into buffer and switches you to that buffer, or switches to

buffer in which the file has previously been read

read-file ^X^R Reads file into buffer thus overwriting buffer contents. If file has already been

read into another buffer, you will be switched to it

view-file ^X^V The same as read-file except the buffer is automatically put into VIEW mode thus preventing any changes from being made

Screen Formatting  MicroEMACS Reference Manual

Chapter 11

Screen Formatting

11.1 Wrapping Text

As we learned in the introduction, EMACS is not a word processor, but an editor. Some simple formatting options are available however, although in most cases they will not affect the appearance of the finished text when it is run through the formatter. We have already encountered WRAP mode which wraps lines longer than a certain length (default is 75 characters). You will recall that WRAP is enabled by entering ^XM and responding to the command line prompt with wrap.

You can also set your own wrap margin with the command ^XF set-fill-column . Notice EMACS responds [Fill column is 1]. Now try typing some text. You’ll notice some very strange things happening — your text wraps at every word!! This effect occurs because the set wrap margin command must be preceded by a numeric argument or EMACS sets it to the first column. Thus any text you type that extends past the first column will wrap at the most convenient line break.

To reset the wrap column to 72 characters, press the <META¿ key and enter 72. EMACS will respond Arg: 72. Now press ^XF. EMACS will respond [Fill column is 72]. Your text will again wrap at the margin you’ve been using up to this point.

11.2 Reformatting Paragraphs

After an intensive editing session, you may find that you have paragraphs containing lines of differing lengths. Although this disparity will not affect the formatted text, aesthetic and technical concerns may make it desirable to have consistent paragraph blocks on the screen. If you are in WRAP mode, you can reformat a paragraph with the command M-Q fillparagraph . This command ”fills” the current paragraph reformatting it so all the lines are filled and wrap logically.

11.3 Changing Case

There may be occasions when you find it necessary to change the case of the text you’ve entered. EMACS allows you to change the case of even large amounts of text with ease. Let’s try and convert a few of the office traditionalists to the joy of word processing. Type in the following text:

Throw away your typewriter and learn to use a word processor. Word processing is relatively easy to learn and will increase your productivity enormously. Enter the Computer Age and find out just how much fun it can be!!

MicroEMACS Reference Manual  Screen Formatting

Let’s give it a little more impact by capitalizing the first four words. The first step is to define the region of text just as you would if you were doing an extensive deletion. Set the mark at the beginning of the paragraph with M-<space¿ set-mark and move the cursor to the space beyond ”typewriter”. Now enter ^X^U case-region-upper. Your text should now look like this:

THROW AWAY YOUR TYPEWRITER and learn to use a word processor. Word processing is relatively easy to learn and will increase your productivity enormously. Enter the Computer Age and find out just how much fun it can be!!

If you want to change the text back to lower case, type ^X^L case-region-lower . You can also capitalize individual words. To capitalize the word fun, position the cursor in front of the word and type M-U case-word-upper . The word is now capitalized. To change it back to lower case, move the cursor back to the beginning of the word and type M-L case-word-lower .

You may also capitalize individual letters in EMACS. The command M-C case-word-capitalize capitalizes the first letter after the point. This command would normally be issued with the cursor positioned in front of the first letter of the word you wish to capitalize. If you issue it in the middle of a word, you can end up with some strAnge looking text.

11.4 Tabs

Unless your formatter is instructed to take screen text literally (as MicroSCRIBE does in the ’verbatim’ environment for example), tabs in EMACS generally affect screen formatting only.

When EMACS is first started, it sets the default tab to every eighth column. As long as you stay with default, every time you press the tab key a tab character, ^I is inserted. This character, like other control characters, is invisible — but it makes a subtle and significant difference to your file and editing.

For example, in default mode, press the tab key and then type the word Test. ‘Test’ appears at the eighth column. Move your cursor to the beginning of the word and delete the backward character. The word doesn’t move back just one character, but flushes to the left margin. The reason for this behavior is easily explained. In tab default, EMACS inserts a ’real’ tab character when you press the tab key. This character is inserted at the default position, but NO SPACES are inserted between the tab character and the margin (or previous tab character). As you will recall, EMACS only recognizes characters (such as spaces or letters) and thus when the tab character is removed, the text beyond the tab is flushed back to the margin or previous tab mark.

This situation changes if you alter the default configuration. The default value may be changed by entering a numeric argument before pressing the tab key. As we saw earlier, pressing the META key and entering a number allows you to specify how EMACS performs a given action. In this case, let’s specify an argument of 10 and hit the tab key.

Now hit the tab key again and type Test. Notice the word now appears at the tenth column. Now move to the beginning of the word and delete the backward character. ”Test” moves back by one character.

EMACS behaves differently in these circumstances because the ^I handle-tab function deals with tabbing in two distinct ways. In default conditions, or if the numeric argument of zero is used, handle-tab inserts a true tab character. If, however, a non-zero numeric argument is specified, handle-tab inserts the correct number of spaces needed to position the cursor at the next specified tab position. It does NOT insert the single tab character and hence any editing functions should take account of the number of spaces between tabbed columns.

The distance which a true tab character moves the cursor can be modified by changing the value of the $hardtab environment variable. Initially set to 8, this will determine how far each tab stop is placed from the previous one. (Use the ^XA set command to set the value of an environment variable).

Many times you would like to take text which has been created using the tab character and change it to use just spaces. The command ^X^D detab-region changes any tabs in the currently selected region into the right number of spaces so the text does not change. This is very useful for times when the file must be printed or transferred to a machine which does not understand tabs.

Also, the inverse command, ^X^E entab-region changes multiple spaces to tabs where possible. This is a good way to shrink the size of large documents, especially with data tables. Both of these commands can take a numeric argument which will be interpreted as the number of lines to en/detab.

Another function, related to those above is provided for by the ^X^T trim-region when invoked will delete any trailing white space in the selected region. A preceding numeric argument will do this for that number of lines.

Chapter 11 Summary

Chapter 11 introduced some of the formatting features of EMACS. Text-wrap, paragraph reformatting, and tabs were discussed in some detail. The commands in the following table were covered in the chapter.

Key Binding  Keystroke  Effect
add-mode/WRAP  ^XM[WRAP] Adds wrap mode to current buffer
delete-mode/WRAP ^X^M[WRAP] Removes wrap mode from current buffer
set-fill-column  ^XF  Sets fill column to given numeric argument
fill-paragraph  M-Q  Logically reformats the current paragraph
case-word-upper  M-U  Text from point to end of the current word is changed to uppercase
case-word-lower  M-L  Text from point to end of the current word is changed to lowercase
case-word-capitalize M-C  First word (or letter) after the point is capitalized
case-region-upper  ^X^U  The current region is uppercased
case-region-lower  ^X^L  The current region is lowercased
handle-tab  ^I  Tab interval is set to the given numeric argument
entab-region  ^X^E  Changes multiple spaces to tabs characters where possible
detab-region  ^X^D  Changes tab characters to the appropriate number of spaces
trim-region  ^X^T  Trims white space from the end of the lines in the current region

36

Access to the Outside World  MicroEMACS Reference Manual

Chapter 12

Access to the Outside World

EMACS has the ability to interface to other programs and the environment of the computer outside of itself. It does this through a series of commands that allow it to talk to the computer’s command processor or shell . Just what this is varies between different computers. Under MSDOS or PCDOS this is the command.com command processor. Under UNIX it is the csh shell. On the Atari ST is can be the Mark Williams MSH or the Beckmeyer shell. In each case, it is the part of the computer’s operating system that is responsible for determining what programs are executed, and when.

The ^X! shell-command command prompts the user for a command line to send out to the shell to execute. This can be very useful for doing file listings and changing the current directory or folder. EMACS gives control to the shell, which executed the command, and then types [END] and waits for the user to type a character before redrawing the screen and resuming editing. If the shell-command command is used from within the macro language, there is no pause.

^X@ pipe-command command allows EMACS to execute a shell command, and if the particular computer allows it, send the results into a buffer which is automatically displayed on the screen. The resulting buffer, called command can be manipulated just like any other editing buffer. Text can be copied out of it or rearranged as needed. This buffer is originally created in VIEW mode, so remember to ^X^Mview<NL¿ in order to change it.

Many computers provide tools which will allow you to filter text, making some modifications to it along the way. A very common tool is the SORT program which accepts a file, sorts it, and prints the result out. The EMACS command, ^X# filter-buffer sends the current buffer through such a filter. Therefore, if you wished to sort the current buffer on a system which supplied a sort filter, you would type ^X#sort<NL¿. You can also create your own filters by writing programs and utilities which read text from the keyboard and display the results. EMACS will use any of these which would normally be available from the current shell.

If you would like to execute another program directly, without the overhead of an intervening shell, you can use the ^X$ execute-program command. It will prompt you for an external program and its arguments and attempt to execute it. Like when EMACS looks for command files, EMACS will look first in the HOME directory, then down the execute PATH, and finally in the current directory for the named program. On some systems, it will automatically tack the proper extension on the file name to indicate it is a program. On some systems that don’t support this function, ^X$ will be equivalent to ^X! shell-command.

Sometimes, you would like to get back to the shell and execute other commands, without losing the current contents of EMACS. The ^XC i-shell command shells out of EMACS, leaving EMACS in the computer and executing another command shell. Most systems would allow you to return to EMACS with the ”exit” command.

On some systems, mainly advanced versions of UNIX, you can direct EMACS to ”go into the background” with the ^XD suspend-emacs command. This places EMACS in the background returning you to the original command shell. EMACS can then be returned to at any time with the ”fg” foreground command.

MicroEMACS Reference Manual  Access to the Outside World

Chapter 12 Summary

Chapter 12 introduced different ways to access the computers shell or command processor from within EMACS. The commands in the following table were covered in the chapter.

Key Binding  Keystroke Effect
execute-program ^X$  Execute an external program directly
filter-buffer  ^X#  Send the current buffer through a shell filter
i-shell  ^XC  Escape to a new shell
pipe-command ^X@  Send the results of an external shell command to a buffer
shell-command ^X!  Execute one shell command
suspend-emacs ^XD  Place EMACS in the background (some UNIX systems only)

Keyboard Macroes  MicroEMACS Reference Manual

Chapter 13

Keyboard Macroes

In many applications, you may need to repeat a series of characters or commands frequently. For example, a paper may require the frequent repetition of a complex formula or a long name. You may also have a series of EMACS commands that you invoke frequently. Keyboard macroes offer a convenient method of recording and repeating these commands.

Imagine, for example, you are writing a scholarly paper on Asplenium platyneuron, the spleenwort fern. Even the dedicated botanist would probably find it a task bordering on the agonizing to type Asplenium platyneuron frequently throughout the paper. An alternative method is ”record” the name in a keyboard macro. Try it yourself.

The command ^X( begin-macro starts recording the all the keystrokes and commands you input. After you’ve typed it, enter Asplenium platyneuron. To stop recording, type ^X) end-macro . EMACS has stored all the keystrokes between the two commands. To repeat the name you’ve stored, just enter ^XE execute-macro , and the name ”Asplenium platyneuron” appears. You can repeat this action as often as you want, and of course as with any EMACS command, you may precede it with a numerical argument to repeat it many times.

Because EMACS records keystrokes, you may freely intermix commands and text. Unfortunately, you can only store one macro at a time. Thus, if you begin to record another macro, the previously defined macro is lost. Be careful to ensure that you’ve finished with one macro before defining another. If you have a series of commands that you would like to ”record” for future use, use the procedure facilities detailed in chapter 14.

Chapter 13 Summary

Chapter 13 covered keyboard macroes. You learned how to record keystrokes and how to repeat the stored sequence.

Key Binding Keystroke Effect
start-macro  ^X(  Starts recording all keyboard input
end-macro  ^X)  Stops recording keystrokes for macro
execute-macro ^XE  Entire sequence of recorded keystrokes is replayed

MicroEMACS Reference Manual  Keyboard Macroes

40

MicroEMACS Procedures  MicroEMACS Reference Manual

Chapter 14

MicroEMACS Procedures

Procedures, or macroes, are programs that are used to customize the editor and to perform complicated editing tasks. They may be stored in files or buffers and may be executed using an appropriate command, or bound to a particular keystroke. Portions of the standard start-up file are implemented via procedures, as well as the built in help system. The M-^E run command causes named procedures to be executed. The execute-file command allows you to execute a procedure stored in a disk file, and the execute-buffer command allows you to execute a procedure stored in a buffer. Procedures are stored for easy execution by executing files that contain the store-procedure command.

In a command file, the store-procedure command takes a string argument which is the name of a procedure to store. These procedures can be executed with the M-^E run command. Also, giving the name of a stored procedure within another procedure will execute that named procedure as if it had been called up with the run command.

Some fairly length examples of MicroEMACS procedures can be seen by examining the standard files that come with EMACS. The emacs.rc file (called .emacsrc under UNIX) is the MicroEMACS command file which is executed when EMACS is normally run. It contains a number of different stored procedures along with the lines to setup and display the Function key window and to call up other procedures and command files using function keys.

There are many different aspects to the language within MicroEMACS. Editor commands are the various commands that manipulate text, buffers, windows, et cetera, within the editor. Directives are commands which control what lines get executed within a macro. Also there are various types of variables. Environmental variables both control and report on different aspects of the editor. User variables hold string values which may be changed and inspected. Buffer variables allow text to be placed into variables. Interactive variable allow the program to prompt the user for information. Functions can be used to manipulate all these variables.

14.1 Constants

All constants and variable contents in EMACS are stored as strings of characters. Numbers are stored digit by digit as characters. This allows EMACS to be ”typeless”, not having different variables types be legal in different contexts. This has the disadvantage of forcing the user to be more careful about the context of the statements variables are placed in, but in turn gives them more flexibility in where they can place variables. Needless to say, this also allows EMACS’s expression evaluator to be both concise and quick.

Wherever statements need to have arguments, it is legal to place constants. A constant is a double quote character, followed by a string of characters, and terminated by another double quote character. To represent various special characters within a constant, the tilde (~) character is used. The character following the tilde is interpreted according to the following table:

MicroEMACS Reference Manual  MicroEMACS Procedures

Sequence Result
~n  EMACS newline character (breaks lines)
~r        ˆ M  carriage return
~l        ˆ J  linefeed
~~        ~  tilde
~b        ˆ H  backspace
~f        ˆ L  formfeed

 ~t        ˆ I  tab

~"  ”  quote

Any character not in the table which follows a tilde will be passed unmodified. This action is similar to the ^Q quotecharacter command available from the keyboard.

EMACS may use different characters for line terminators on different computers. The ~n combination will always get the proper line terminating sequence for the current system.

The double quotes around constants are not needed if the constant contains no internal white space and it also does not happen to meet the rules for any other EMACS commands, directives, variables, or functions. This is reasonable useful for numeric constants.

14.2 Variables

Variables in MicroEMACS procedures can be used to return values within expressions, as repeat counts to editing commands, or as text to be inserted into buffers and messages. The value of these variables is set using the set ^XA command. For example, to set the current fill column to 64 characters, the following macro line would be used:

     set $fillcol 64

or to have the contents of %name inserted at the point in the current buffer, the command to use would be:

     insert-string %name

14.2.1 Environmental Variables

”What good is a quote if you can’t change it?”

These variables are used to change different aspects of the way the editor works. Also they will return the current settings if used as part of an expression. All environmental variable names begin with a dollar sign ($) and are in lower case.

$acount The countdown of inserted characters until the next save-file.

$asave The number of inserted characters between automatic file-saves in ASAVE mode.

$bufhook The function named in this variable is run when a buffer is entered. It can be used to implement modes which are specific to a particular file or file type.

$cbflags Current buffer attribute flags (See appendix G for details).

$cbufname Name of the current buffer.

$cfname File name of the current buffer.

$cmdhook Name of function to run before accepting a command. This is by default set to nop.

$cmode Integer containing the mode of the current buffer. (See Appendix F for values).

$curchar Ascii value of the character currently at the point.

$curcol Current column of point in current buffer. $curline Current line of point in current buffer. $curwidth Number of columns used currently. $curwind Current window number. $cwline Current display line in current window. $debug Flag to trigger macro debugging. $deskcolor Color to use for current desktop, default to BLACK. $diagflag If set to TRUE, diagonal dragging of text and mode lines is enabled. If FALSE, text and modelines can only

be dragged horizontally or vertically at one time.

$discmd Controls the echoing of command prompts. Default is TRUE. $disinp Controls the echoing of input at the command prompts. Default is TRUE. $disphigh If set to TRUE, high-bit characters (single byte characters that are greater than 127 in value) will be displayed

in a pseudo-control format. The characters ‘^!’ will lead off the sequence, followed by the character stripped of its high bit. Default is FALSE.

$exbhook This variable holds the name of a function or macro which is run whenever you are switching out of a buffer. $fcol The current line position being displayed in the first column of the current window. $fillcol Current fill column. $flicker Flicker Flag set to TRUE if IBM CGA set to FALSE for most others. $fmtlead lists all formatter command leadin characters. Lines beginning with these characters will be considered the

beginning of paragraphs.

$gflags Global flags controlling some EMACS internal functions (See appendix G for details). $gmode Global mode flags. (See Appendix F for values). $hardtab Number of spaces between hard tab stops. Normally 8, this can be used to change indentation only within

the editor.

$hjump The number in here tells EMACS how many columns to scroll the screen horizontally when a horizontal

scroll is required.

$hilight Indicates to the display driver that the region between the point and the indicated mark should be hilighted.

A value of 255 indicates no hilighting.

$hscroll This flag determines if EMACS will scroll the entire current window horizontally, or just the current line. The default value, TRUE, results in the entire current window being shifted left and right when the cursor goes off the edge of the screen.

$kill This contains the first 127 characters currently in the kill buffer and can be used to set the contents of the kill

buffer.

$isterm This is the character used to terminate incremental search string inputs. The default for this is the last key

bound to meta-prefix.

$language [READ ONLY] Contains the name of the language which the current EMACS’s message will display. (Cur-

rently EMACS is available in English, French, Spanish, Latin, Portuguese, Dutch, German, and Pig Latin).

$lastkey [READ ONLY] Last keyboard character typed. $lastmesg [READ ONLY] Contains the text of the last message which emacs wrote on the command line.

$line The current line in the current buffer can be retrieved and set with this environment variable. $lterm Character(s) to write as a line terminator when writing a file to disk. Default is null, which causes a \n

character to be written. Not all operating systems support this.

$lwidth [READ ONLY] Returns the number of characters in the current line. $match [READ ONLY] Last string matched in a search. $modeflag Determines if mode lines are currently displayed. $msflag If TRUE, the mouse (if present) is active. If FALSE, no mouse cursor is displayed, and no mouse actions are

taken.

$numwind The number of windows displayed. $oldcrypt Use the old method of encryption (which had a bug in it). Default is FALSE. This variable was dropped from

EMACS with version 3.12.

$orgcol The desktop column position of current screen. $orgrow The desktop row position of current screen. $os The system name under which you run EMACS. $overlap Controls number of overlapped lines during paging. Default value is 2. $pagelen The number of screen lines used currently. $palette A string used to control the palette register settings on graphics versions. The usual form consists of groups

of three octal digits setting the red, green, and blue levels.

$paralead A string containing all paragraph start characters. $pending [READ ONLY] A flag used to determine if there are user keystrokes waiting to be processed. $popflag Use pop-up windows. Default is TRUE. $posflag Display the line and column position on the modeline. Default is FALSE. $progname [READ ONLY] Always contains the string ”MicroEMACS” for standard MicroEMACS. Could be something

else if EMACS is incorporated as part of someone else’s program.

$ram The amount of remaining memory if MicroEMACS was compiled with RAMSIZE set. A debugging tool. $readhook This variable holds the name of a function to execute whenever a file is read into EMACS. Normally, using

the standard emacs.rc file, this is bound to a function which places EMACS into CMODE if the extension of the file read is .c or .h.

$region Contains the string of the current region. It will truncate at the stringsize limit, 255. $replace The current replace pattern used in replace commands. $rval This contains the return value from the last subprocess which was invoked from EMACS. $scrname The current screen name. $search The current search pattern used in search and replace commands. $searchpnt Set the placement of the cursor on a successful search match. $searchpnt = 0 (the default), causes the cursor

to be placed at the end of the matched text on forward searches, and at the beginning of the text on reverse searches. $searchpnt = 1 causes the cursor to be placed at the the beginning of the matched text regardless of the search direction, while $searchpnt = 2 causes the cursor to be placed at the end.

$seed Integer seed of the random number generator.

$softtab Number of spaces inserted by EMACS when the handle-tab command (which is normally bound to the TAB key) is invoked.

$sres Current screen resolution (CGA, MONO, EGA or VGA on the IBM-PC driver. LOW, MEDIUM, HIGH or DENSE on the Atari ST1040, NORMAL on most others).

$ssave A variable which flags EMACS’s method of saving files. If set to TRUE, EMACS will write all files out to a temporary file, delete the original, then rename the temporary to the old file name. The default value of this is TRUE.

$sscroll When set to TRUE, EMACS will smoothly scroll windows one line at a time when cursoring off the ends of the current window. Default is FALSE.

$status [READ ONLY] Status of the success of the last command (TRUE or FALSE). This is usually used with !force to check on the success of a search, or a file operation.

$sterm This is the character used to terminate search string inputs. The default for this is the last key bound to meta-prefix.

$target Current target for line moves (setting this fools EMACS into believing the last command was a line move).

$time [READ ONLY] Contains a string corresponding to the current date and time. Usually this is in a form similar to ”Mon May 09 10:10:58 1988”. Not all operating systems will support this.

$timeflag Flag to determine if the time of day is displayed on the modeline. Default is FALSE. The time is updated only after a keystroke.

$tpause Controls the length of the pause to display a matched fence when the current buffer is in CMODE and a close fence has been typed.

$version [READ ONLY] Contains the current MicroEMACS version number.

$wchars When set, MicroEMACS uses the characters listed in it to determine if it is in a word or not. If it is not set (the default), the characters it uses are the upper and lower case letters, and the underscore.

$wline Number of display lines in current window.

$wraphook This variable contains the name of an EMACS function which is executed when a buffer is in WRAP mode and it is time to wrap. By default this is bound to wrap-word.

$writehook This variable contains the name of an EMACS function or macro which is invoked whenever EMACS attempts to write a file out to disk. This is executed before the file is written, allowing you to process a file on the way out.

$xpos The column the mouse was at the last mouse button press.

$yankflag Controls the placement of the cursor after a yank command or an insert. When $yankflag is FALSE (the default), the cursor is placed at the end of the yanked or inserted text. When it is TRUE, the cursor remains at the start of the text.

$ypos The line which the mouse was on during the last mouse button press.

14.2.2 User variables

User variables allow you to store strings and manipulate them. These strings can be pieces of text, numbers (in text form), or the logical values TRUE and FALSE. These variables can be combined, tested, inserted into buffers, and otherwise used to control the way your macroes execute. At the moment, up to 512 user variables may be in use in one editing session. All users variable names must begin with a percent sign (%) and may contain any printing characters. Only the first 10 characters are significant (IE differences beyond the tenth character are ignored). Most operators will truncate strings to a length of 128 characters.

14.2.3 Buffer Variables

Buffer variables are special in that they can only be queried and cannot be set. What buffer variables are is a way to take text from a buffer and place it in a variable. For example, if I have a buffer by the name of RIGEL2, and it contains the text:

              Richmond
              Lafayette
              <*>loomington             (where <*> is the current point)
              Indianapolis
              Gary
              =* MicroEMACS 3.12 (WRAP) == rigel2 == File: /data/rigel2.txt =====

and within a command I reference #rigel2, like:

     insert-string #rigel2

MicroEMACS would start at the current point in the RIGEL2 buffer and grab all the text up to the end of that line and pass that back. Then it would advance the point to the beginning of the next line. Thus, after our last command executes, the string ”Bloomington” gets inserted into the current buffer, and the buffer RIGEL2 now looks like this:

              Richmond
              Lafayette
              Bloomington
              <*>Indianapolis          (where <*> is the current point)
              Gary
              =* MicroEMACS 3.12 (WRAP) == rigel2 == File: /data/rigel2.txt =====

as you have probably noticed, a buffer variable consists of the buffer name, preceded by a pound sign (#).

14.2.4 Interactive variables

Interactive variables are actually a method to prompt the user for a string. This is done by using an at sign (@) followed either with a quoted string, or a variable containing a string. The string is then placed on the bottom line, and the editor waits for the user to type in a string. Then the string typed in by the user is returned as the value of the interactive variable. For example:

              set %quest "What file? "
              find-file @%quest

will ask the user for a file name, and then attempt to find it. Note also that complex expressions can be built up with these operators, such as:

     set %default "file1"
     @&cat &cat "File to decode[" %default "]: "

which prompts the user with the string:

     File to decode[file1]:

14.3 Functions

Functions can be used to act on variables in various ways. Functions can have one, two, or three arguments. These arguments will always be placed after the function on the current command line. For example, if we wanted to increase the current fill column by two, using emacs’s set (^XA) command, we would write:

        set $fillcol &add $fillcol 2
         \       \         \        \      \___second operand
          \       \         \        \________first operand
           \       \         \_______________function to execute
             \       \_______________________variable to set
              \_____________________________set (^XA) command

Function names always begin with the ampersand (&) character, and are only significant to the first three characters after the ampersand. Functions will normal expect one of three types of arguments, and will automatically convert types when needed. Different argument types include:

<num> an ascii string of digits which is interpreted as a numeric value. Any string which does not start with a digit or a minus sign (−) will be considered zero.

<str> An arbitrary string of characters. At the moment, strings are limited to 128 characters in length.

<log> A logical value consisting of the string ”TRUE” or ”FALSE”. Numeric strings will also evaluate to ”FALSE” if they are equal to zero, and ”TRUE” if they are non-zero. Arbitrary text strings will have the value of ”FALSE”.

A list of the currently available functions follows. Functions are always used in lower case, the uppercase letters in the function table are the short form of the function (IE &div for &divide).

Numeric Functions: (returns <num>)

&ADD  <num> <num> Add two numbers
&SUB  <num> <num> Subtract the second number from the first
&TIMes  <num> <num> Multiply two numbers
&DIVide <num> <num> Divide the first number by the second giving an integer result
&MOD  <num> <num> Return the reminder of dividing the first number by the second
&NEGate <neg>  Multiply the arg by -1
&LENgth <str>  Return length of string
&SINdex <str1> <str2> Finds the position of <str2> within <str1>. Returns zero if not found.
&ASCii  <str>  Return the ascii code of the first character in <str>
&RND  <num>  Return a random integer between 1 and <num>
&ABS  <num>  Return the absolute value of <num>
&BANd  <num> <num> Bitwise AND function
&BOR  <num> <num> Bitwise OR function
&BXOr  <num> <num> Bitwise XOR function
&BNOt  <num>  Bitwise NOT function

String manipulation functions: (returns <str>)

&CAT  <str> <str>  Concatenate the two strings to form one
&LEFt  <str> <num>  Return the <num> leftmost characters from <str>
&RIGht  <str> <num>  Return the <num> rightmost characters from <str>
&MID  <str> <num1> <num2> Starting from <num1> position in <str>, return <num2> characters.
&REVerse <str>  Return a string with reversed-ordered characters
&UPPer  <str>  Uppercase <str>
&LOWer  <str>  Lowercase <str>
&CHR  <num>  Return a string with the character represented by ascii code <num>

&GTC Return a string of characters containing a EMACS command input from

the user

&GTK  Return a string containing a single keystroke from the user

&ENV <str> If the operating system is capable, this returns the environment string

associated with <str>

&BIND <str> Return the function name bound to the keystroke <str>
&XLATE <str1> <str2> <str3>

&FINd <str> Find the named file <str> along the path and return its full file specifi-

cation or an empty string if none exists

&TRIM  <str>  Trim the trailing whitespace from a string

Logical Testing functions: (returns <log>)

&NOT  <log>  Return the opposite logical value
&AND  <log1> <log2>  Return TRUE if BOTH logical arguments are TRUE
&OR  <log1> <log2>  Return TRUE if either argument is TRUE
&EQUal  <num> <num>  If <num> and <num> are numerically equal, return TRUE
&LESs  <num1> <num2> If <num1> is less than <num2>, return TRUE.
&GREater <num1> <num2> If <num1> is greater than <num2>, return TRUE.
&SEQual  <str1> <str2>  If the two strings are the same, return TRUE.
&SLEss  <str1> <str2>  If <str1> is less alphabetically than <str2>, return TRUE.
&SGReater <str1> <str2>  If <str1> is alphabetically greater than <str2>, return TRUE.
&EXIst  <str>  Does the named file <str> exist?
&ISNum  <num>  Is the given argument a legitimate number?

Special Functions:

&GROup <num>  Return group <num> as set by a MAGIC mode search.

&SUPper <str1> <str2> Translate the first char in <str1> to the first char in <str2> when up-

percasing.

&SLOwer <str1> <str2> Translate the first char in <str1> to the first char in <str2> when low-

ercasing.

&INDirect <str>  Evaluate <str> as a variable.

This last function deserves more explanation. The &IND function evaluates its argument, takes the resulting string, and then uses it as a variable name. For example, given the following code sequence:

             ; set up reference table
             set %one         "elephant"
             set %two         "giraffe"
             set %three       "donkey"
             set %index "%two"
             insert-string &ind %index

the string ”giraffe” would have been inserted at the point in the current buffer. This indirection can be safely nested up to about 10 levels.

14.4 Directives

Directives are commands which only operate within an executing procedure, IE they do not make sense as a single command. As such, they cannot be called up singly or bound to keystroke. Used within command files, they control what lines are executed and in what order.

Directives always start with the exclamation mark (!) character and must be the first non-white space placed on a line. Directives executed interactively (via the execute-command-line command) will be ignored.

14.4.1 !ENDM Directive

This directive is used to terminate a procedure or macro being stored. For example, if a file is being executed contains the text:

              ;         Read in a file in view mode, and make the window red
              store-procedure get-red-viewed-file
                       find-file @"File to view: "
                       add-mode "view"
                       add-mode "red"
              !endm
              print "[Consult procedure has been loaded]"

only the lines between the store-macro command and the !ENDM directive are stored in procedure get-red-viewd-file. Both named procedures and numbered macroes (via the store-macro command) should be terminated with this directive.

14.4.2 !FORCE Directive

When MicroEMACS executes a procedure, if any command fails, the procedure is terminated at that point. If a line is preceded by a !FORCE directive, execution continues whether the command succeeds or not. For example:

              ;         Merge the top two windows
              save-window                  ;remember what window we are at
              1 next-window                  ;go to the top window
              delete-window                  ;merge it with the second window
              !force restore-window         ;This will continue regardless
              add-mode "red"

Often this is used together with the $status environment variable to test if a command succeeded. For example:

              set %seekstring @"String to Find: "
              !force search-forward %seekstring
              !if &seq $status TRUE
                       print "Your string is Found"
              !else
                       print "No such STRING!"
              !endif

14.4.3 !IF, !ELSE, and !ENDIF Directives

This directive allows statements only to be executed if a condition specified in the directive is met. Every line following the !IF directive, until the first !ELSE or !ENDIF directive, is only executed if the expression following the !IF directive evaluates to a TRUE value. For example, the following commands creates the portion of a text file automatically. (yes believe me, this will be easier to understand then that last explanation....)

!if &sequal %curplace "timespace vortex"
insert-string "First, rematerialize~n"
!endif
!if &sequal %planet "earth" ;If we have landed on earth...
!if &sequal %time "late 20th century" ;and we are then
write-message "Contact U.N.I.T."
!else
insert-string "Investigate the situation....~n"
insert-string "(SAY 'stay here Sara')~n"
!endif
!else
set %conditions @"Atmosphere conditions outside?  "
!if &sequal %conditions "safe"
insert-string &cat "Go outside......" "~n"
insert-string "lock the door~n"
!else
insert-string "Dematerialize..try somewhen else"
newline
!endif
!endif

14.4.4 !GOTO Directive

Flow can be controlled within a MicroEMACS procedure using the !GOTO directive. It takes as an argument a label. A label consists of a line starting with an asterisk (*) and then an alphanumeric label. Only labels in the currently executing procedure can be jumped to, and trying to jump to a non-existing label terminates execution of a procedure. For example:

    ;Create a block of DATA statements for a BASIC program
              insert-string "1000 DATA "
              set %linenum 1000
    *nxtin
              update-screen                  ;make sure we see the changes
              set %data @"Next number: "
              !if &equal %data 0
                       !goto finish
              !endif
              !if &greater $curcol 60
                       2 delete-previous-character
                       newline
                       set %linenum &add %linenum 10
                       insert-string &cat %linenum " DATA "
              !endif
              insert-string &cat %data ", "
              !goto nxtin

         *finish
                  2 delete-previous-character
                  newline

14.4.5 !WHILE and !ENDWHILE Directives

This directive allows you to set up repetitive tasks easily and efficiently. If a group of statements need to be executed while a certain condition is true, enclose them with a while loop. For example,

              !while &less $curcol 70
                       insert-string &cat &cat "[" #stuff "]"
              !endwhile

places items from buffer ”item” in the current line until the cursor is at or past column 70. While loops may be nested and can contain and be the targets of !GOTOs with no ill effects. Using a while loop to enclose a repeated task will run much faster than the corresponding construct using !IFs.

14.4.6 !BREAK Directive

This lets you abort out of the most executing currently inner while loop, regardless of the condition. It is often used to abort processing for error conditions. For example:

     ;         Read in files and substitute "begining" with "beginning"
              set %filename #list
              !while &not &seq %filename "<end>"
     !force                  find-file %filename
                       !if &seq $status FALSE
                                write-message "[File read error]"
                                !break
                       !endif
                       beginning-of-file
                       replace-string "begining" "beginning"
                       save-file
                       set %filename #list
              !endwhile

This while loop will process files until the list is exhausted or there is an error while reading a file.

14.4.7 !RETURN Directive

The !RETURN Directive causes the current procedure to exit, either returning to the caller (if any) or to interactive mode. For example:

     ; Check the monitor type and set %mtyp
     !if &sres "CGA"
     set %mtyp 1
     !return
     !else

set %mtyp 2
!endif
insert-string "You are on a MONOCHROME machine!~n"

Debugging MicroEMACS Procedures  MicroEMACS Reference Manual

Chapter 15

Debugging MicroEMACS Procedures

When developing new procedures, it is very convenient to be able to trace their execution to discover errors. The $debug environment variable enables procedure debugging. While this variable is TRUE, emacs will stop at each line it intends to execute and allow you to view it, and issue a number of different commands to help determine how the procedure is executing.

For example, we will step through the procedure which toggles the function key window off. The first thing to do, is to set $debug, using the ^XA set command. Type ^XA and emacs will prompt you on the command line with Variable to set: . Type in $debug and press the enter key. Emacs will then ask Value: . Type in TRUE (in capital letters) and press the enter key.

While macro debugging is enabled (as it is now) emacs will report each time a variable is assigned a value, by displaying the variable and its value on the command line. Right now,

     ((($debug <- TRUE)))

appears on the command line to tell you that $debug now has been assigned the value of TRUE. Press the space bar to continue.

Now, lets try to debug a macro. Press function key 5 which normally toggles the function key window. The first thing that appears is:

     <<<[Macro 01]:!if %rcfkeys>>>

At this point, emacs is waiting for a command. It is prepared to see if the user variable %rcfkeys is TRUE, and execute some lines if they are. Suppose we want to see the value of this variable, type the letter e to evaluate an expression. Emacs will prompt with EXP: . Type %rcfkeys followed by the enter key. Emacs should then respond with TRUE to indicate that the function key window is currently on screen.

Press the space bar to allow the !if directive to execute. Emacs will decide that it is TRUE, and then display the next command to execute.

     <<<[Macro 01]:!goto rcfoff>>>

Notice emacs tells us what procedure we are currently executing (in this case, the macro bound to execute-macro-1). Press the space bar again to execute the !goto directive.

     <<<[Macro 01]:save-window>>>

Emacs is saving the position of the current window so that it can attempt to return to it after it has brought up the function key window.

[...THIS CHAPTER IS NOT FINISHED...]

MicroEMACS Reference Manual  Debugging MicroEMACS Procedures

54

Key Bindings, What they are and why  MicroEMACS Reference Manual

Chapter 16

Key Bindings, What they are and why

One of the features which makes MicroEMACS very adaptable is its ability to use different keystrokes to execute different commands. The process of changing the particular command a key invokes is called rebinding. This allows us to make the editor look like other popular editors and programs.

Each command in MicroEMACS has a name which is used for binding purposes. For example, the command to move the cursor down one page is called next-line and is normally bound to the ^N key. If you decided that you also wanted to use the ^D key to move the cursor down one line, you would use the M-K bind-to-key command. EMACS would respond with : bind-to-key on the command line and allow you to type in a command name. Then type in the name of the command you want to change, in this case next-line, followed by the <NL> key. EMACS will then wait for you to type in the keys you want to activate the named function. Type a single ^D. From now on, typing ^D will cause EMACS to move down one line, rather than its original function of deleting characters.

To find out the name of a command, consult the list of valid EMACS commands in Appendix B. Also, you can use the ^X? describe-key command to look up the name of a command. Type ^X? and then the key to use that command, and EMACS will show you the name of the command.

After you have experimented with changing your key bindings, you may decide that you want to change some bindings permanently. To have EMACS rebind keys to your pleasure each time you start EMACS, you can add statements to the end of your startup file (emacs.rc or .emacsrc depending on the system). For example,

bind-to-key next-line ^D

Notice, that control D character in the startup file is represented visibly as an uparrow key followed by a capital D. To know how to represent any keys you want to bind, use the describe-key command on the key, and use the sequence that is displayed.

bind-to-key split-current-window FN1

This example would make function key 1 activate the command that splits the current window in two.

EMACS will let you define a large number of keys, but will report Binding table FULL! when it runs out of space to bind keys. Normally EMACS will allow up to 512 key bindings (including approx. 300 originally bound keys).

If you want to get a current listing of all the commands and the keys bound to them, use the describe-bindings command. Notice, that this command is not bound to any keys!

There are some key bindings that cannot be made without special precautions. Alternative bindings for ^X, META, ^G, and ^U (which bind respectively to ctlx-prefix, meta-prefix, abort-command, and universal-argument) must be made before re-binding ^X, META, ^G, or ^U. The reason is to protect the innocent user from losing the prefix and other commands inadvertantly.

MicroEMACS Reference Manual  Key Bindings, What they are and why

56

MicroEMACS Command Line Switches and Startup Files  MicroEMACS Reference Manual

Appendix A

MicroEMACS Command Line Switches and Startup Files

When EMACS first executes, it always searches for a file, called .emacsrc under most UNIX systems or emacs.rc on most other systems which it will execute as EMACS macroes before it reads in the named source files. This file normally contains EMACS macroes to bind the function keys to useful functions and load various useful macroes. The contents of this file will probably vary from system to system and can be modified by the user as desired.

When searching for this file, EMACS looks for it in this order. First, it attempts to find a definition for ‘HOME’ in the environment. It will look in that directory first. Then it searches all the directories listed in the ‘PATH’ environment variable. Then it looks through a list of predefined standard directories which vary from system to system. Finally, failing all of these, it looks in the current directory. This is also the same method EMACS uses to look up any files to execute, and to find it’s help file emacs.hlp.

On computers that call up EMACS via a command line process, such as MSDOS and UNIX, there are different things that can be added to the command line to control the way EMACS operates. These can be switches, which are a dash (–) followed by a letter, and possible other parameters, or a startup file specifier, which is an at sign @ followed by a file name.

@<file> This causes the named file to be executed instead of the standard emacs.rc file before emacs reads in any other files. More than one of these can be placed on the command line, and they will be executed in the order that they appear.

–C The following source files on the command line can be changed (as opposed to being in VIEW mode). This is mainly used to cancel the effects of the –v switch used previously in the same command line.

–E This flag causes emacs to automatically run the startup file error.cmd instead of emacs.rc. This is used by various C compilers for error processing (for example, Mark Williams C).

–G<num> Upon entering EMACS, position the cursor at the <num> line of the first file.

+<num> Placing a +¡num¿ on the command line is equivalant to -g¡n¿, allowing the point to be placed on any line of the first file listed on the command line.

–P<num> Upon entering EMACS, position the cursor at the <num> character of the first file. This can be used in conjunction with the –G flag to position the point anywhere within that file.

–I<var> <value> Initialize an EMACS variable with <value>. This can be useful to force EMACS to start in a particular mode. (For example, invoke EMACS with emacs –i$sres VGA foo.bar to edit file foo.bar in VGA 50 line mode on an IBM-PC).

–K<key> This key tells emacs to place the source files in CRYPT mode and read it in using <key> as the encryption key. If no key is listed immediately after the –K switch, EMACS will prompt for a key, and not echo it as it is typed.

MicroEMACS Reference Manual  MicroEMACS Command Line Switches and Startup Files

–R This places EMACS in ”restricted mode” where any commands allowing the user to read or write any files other than the ones listed on the command line are disabled. Also all commands allowing the user access to the operating system are disabled. This makes EMACS very useful as a ”safe” environment for use within other applications and especially used as a remote editor for a BBS or electronic bulletin board system.

–S<string> After EMACS is started, it automatically searches for <string> in the first source file.

–V This tells EMACS that all the following sources files on the command line should be in VIEW mode to prevent any changes being made to them.

Command Completion  MicroEMACS Reference Manual

Appendix B

Command Completion

Some versions of MicroEMACS will allow you to abbreviate buffer names, command names and file names as you enter them. To use this, type in the first few characters of the name you wish, and then hit either the space bar, the META key or the TAB key. MicroEMACS will then attempt to look at the list of all the availible names and if there is only one which will fit, it will choose that name. If there are several names that quailify, as many characters as are common to ALL of them will be entered. If there are no possible matches, the bell will ring to indicate MicroEMACS can not complete the command.

For example, if you have several files in your current directory with the following names:

              prog1.c
              prog1.obj
              prog1.exe
              prog1.doc
              program.one
              project.one
              test.c
              tes

and you enter the ^X^F find-file command, if you type p and then hit the space bar, EMACS will respond by typing the r that is common to all the above file names begining with p. If you then type ogr and hit the tab key, EMACS will respond with am.one and automatically hit the enter key for you.

If you were to instead type an a and hit the space bar, EMACS will beep, informing you that there is no possible match.

If you type a te and hit the space bar, EMACS will then type the following s, but it will not automatically enter it because it is possible you mean to get to the test.c file.

Buffer name, and command name completion is available in all versions of MicroEMACS. File name completion is available on UNIX BSD4.3, the Atari ST, the AMIGA and under MSDOS.

The ’ ?’ or an ambiguous completion causes a pop-up buffer containing the names of all files, commands or buffers that could match a partial completion. Very handy when searching through file structures.

MicroEMACS Reference Manual  Command Completion

60

MicroEMACS Commands  MicroEMACS Reference Manual

Appendix C

MicroEMACS Commands

Below is a complete list of the commands in EMACS, the keys normally used to do the command, and what the command does. Remember, on some computers there may also be additional ways of using a command (cursor keys and special function keys for example).

Command  Binding  Meaning

abort-command

                               ^G

This allows the user to abort out of any command that is waiting for input

add-global-mode M-M Add a global mode for all new buffers
add-mode                       ^XM  Add a mode to the current buffer
append-file                    ^X^A  Write a buffer to the end of a file
apropos M-A List out commands whose name contains the string specified
backward-character             ^B  Move one character to the left
begin-macro                    ^X(  Begin recording a keyboard macro
beginning-of-file M-< Move to the beginning of the file in the current buffer
beginning-of-line              ^A  Move to the beginning of the current line
bind-to-key  M-K  Bind a key to a function

buffer-position

^X= List the position of the cursor in the current window on the command line

case-region-lower              ^X^L  Make a marked region all lower case
case-region-upper              ^X^U  Make a marked region all upper case
case-word-capitalize  M-C  Capitalize the following word
case-word-lower  M-L  Lower case the following word
case-word-upper  M-U  Upper case the following word
change-file-name               ^XN  Change the name of the file in the current buffer
change-screen-size  (none)  Change the number of lines of the screen currently being used
change-screen-width  (none)  Change the number of columns of the screen currently being used
clear-and-redraw               ^L  Clear the physical screen and redraw it
clear-message-line  (none)  Clear the command line
copy-region  M-W  Copy the currently marked region into the kill buffer

count-words

M-^C Count how many words, lines and characters are in the current marked region

ctlx-prefix                    ^X  Change the key used as the ^X prefix
cycle-ring                     ^X-Y  Move the current position in the ring

MicroEMACS Reference Manual  MicroEMACS Commands

Command  Binding  Meaning
cycle-screens  A-C  Bring the rearmost screen to front
delete-blank-lines             ^X^O  Delete all blank lines around the cursor
delete-buffer                  ^XK  Delete a buffer which is not being currently displayed in a window
delete-kill-ring  M-^Y  Dump all the killed text in the ring
delete-global-mode  M-^M  Turn off a global mode
delete-mode                    ^X^M  Turn off a mode in the current buffer
delete-next-character          ^D  Delete the character following the cursor
delete-next-word M-D Delete the word following the cursor
delete-other-windows           ^X1  Make the current window cover the entire screen
delete-previous-character      ^H  Delete the character to the left of the cursor
delete-previous-word  M-^H  Delete the word to the left of the cursor
delete-screen  A-D  Delete a screen
delete-window                  ^X0  Remove the current window from the screen
describe-bindings  (none)  Make a list of all legal commands
describe-functions  (none)  Make a list of all legal functions
describe-variables  (none)  Make a list of all environment and user variables
describe-key                   ^X?  Describe what command is bound to a keystroke sequence
detab-region                   ^X^D  Change all tabs in a region to the equivalent spaces
display                        ^XG  Prompts the user for a variable and displays its current value

dump-variables

(none)

Places into a buffer the current values of all environment and user variables

end-macro                      ^X)  stop recording a keyboard macro
end-of-file M-> Move cursor to the end of the current buffer
end-of-line                    ^E  Move to the end of the current line
end-of-word (none) Move the point just past the end of the current word
entab-region                   ^X^E  Change multiple spaces to tabs where possible

exchange-point-and-mark

                               ^X^X

Move cursor to the last marked spot, make the original position be marked

execute-buffer  (none)  Execute a buffer as a macro
execute-command-line  M-^X  Execute a line typed on the command line as a macro command
execute-file  (none)  Execute a file as a macro
execute-macro                  ^XE  Execute the keyboard macro (play back the recorded keystrokes)

execute-macro-<n>

(none)

Execute numbered macro <N> where <N> is an integer from 1 to 40

execute-named-command  M-X  Execute a command by name
execute-procedure  M-^E  Execute a procedure by name
execute-program                ^X$  Execute a program directly (not through an intervening shell)

exit-emacs

                               ^X^C

Exit EMACS. If there are unwritten, changed buffers EMACS will ask to confirm

fill-paragraph M-Q Fill the current paragraph
filter-buffer                  ^X#  Filter the current buffer through an external filter
find-file                      ^X^F  Find a file to edit in the current window
find-screen A-F Bring the named screen to front, creating it if needed
forward-character              ^F  Move cursor one character to the right
goto-line  M-G  Goto a numbered line
goto-mark  M-^G  Goto a numbered mark
goto-matching-fence  M-^F  Goto the matching fence

Command  Binding  Meaning
grow-window                    ^X^Make the current window larger
handle-tab                     ^I  Insert a tab or set tab stops
hunt-forward  A-S  Hunt for the next match of the last search string
hunt-backward  A-R  Hunt for the last match of the last search string
help  M-?  Read EMACS.HLP into a buffer and display it
i-shell                        ^XC  Shell up to a new command processor
incremental-search             ^XS  Search for a string, incrementally
indent-region M-( Indent the current region one tab
insert-file                    ^X^I  insert a file at the cursor in the current file
insert-space                   ^C  Insert a space to the right of the cursor
insert-string  (none)  Insert a string at the cursor
kill-paragraph  M-^W  Delete the current paragraph
kill-region                    ^W  Delete the current marked region, moving it to the kill buffer
kill-to-end-of-line            ^K  Delete the rest of the current line
label-function-key (none) Set the text on a function key label (HP150 only)
list-buffers                   ^X^B  List all existing buffers
list-screens  A-B  List all existing screens
macro-to-key  M-^K  Bind a key to a macro
meta-prefix  <ESC>  Key used to precede all META commands

mouse-move-down  MSa
mouse-move-up  MSb
mouse-resize-screen  MS1
mouse-region-down  MSe
mouse-region-up  MSf

move-window-down               ^X^N  Move all the lines in the current window down
move-window-up                 ^X^P  Move all the lines in the current window up
name-buffer M-^N Change the name of the current buffer
narrow-to-region               ^X<  hides all text not in the current region
newline                        ^M  Insert a <NL> at the cursor

newline-and-indent

                               ^J

Insert a <NL> at the cursor and indent the new line the same as the preceding line

next-buffer                    ^XX  Bring the next buffer in the list into the current window
next-line                      ^N  Move the cursor down one line
next-page                      ^V  Move the cursor down one page
next-paragraph M-N Move cursor to the next paragraph
next-window                    ^XO  Move cursor to the next window
next-word  M-F  Move cursor to the beginning of the next word
nop  (none)  Does nothing
open-line                      ^O  Open a line at the cursor
overwrite-string (none) Overwrite a string at the cursor
pipe-command                   ^X@  Execute an external command and place its output in a buffer
pop-buffer (none) Display a buffer temporarily, paging
previous-line                  ^P  Move cursor up one line
previous-page                  ^Z  Move cursor up one page
previous-paragraph M-P Move back one paragraph
previous-window                ^XP  Move the cursor to the last window
previous-word  M-B  Move the cursor to the beginning of the word to the left of the cursor

Command  Binding  Meaning
print  (none)  Display a string on the command line (a synonym to write-message)

query-replace-string

M-^R Replace all of one string with another string, interactively querying the user

quick-exit M-Z Exit EMACS, writing out all changed buffers
quote-character                ^Q  Insert the next character literally
read-file                      ^X^R  Read a file into the current buffer
redraw-display  M-^L  Redraw the display, centering the current line
remove-mark  (none)  Remove a numbered mark

replace-string

M-R Replace all occurrences of one string with another string from the cursor to the end of the buffer

resize-window                  ^XW  Change the number of lines in the current window
restore-window (none) Move cursor to the last saved window
reverse-incremental-search     ^XR  Search backwards, incrementally
run M-^E Execute a named procedure
save-file                      ^X^S  Save the current buffer if it is changed
save-window  (none)  Remember current window (to restore later)
scroll-next-up  M-^Z  Scroll the next window up
scroll-next-down  M-^V  Scroll the next window down
search-forward                 ^S  Search for a string
search-reverse                 ^R  Search backwards for a string
select-buffer                  ^XB  Select a buffer to display in the current window
set                            ^XA  Set a variable to a value
set-encryption-key M-E Set the encryption key of the current buffer
set-fill-column                ^XF  Set the current fill column
set-mark  Set the mark
shell-command                  ^X!  Causes an external shell to execute a command
show-files (none) Pop up a list of files from the specified directory
shrink-window                  ^X^Z  Make the current window smaller
source (none) Execute a file as a macro
split-current-window           ^X2  Split the current window in two
store-macro  (none)  Store the following macro lines to a numbered macro
store-procedure  (none)  Store the following macro lines to a named procedure

suspend-emacs

^XD Return you to the original command shell, EMACS being returned to with the ”fg” foreground command (Unix and VMS)

transpose-characters           ^T  Transpose the character at the cursor with the character to the left
trim-region                    ^X^T  Trim any trailing white space from a region
unbind-key  M-^K  Unbind a key from a function
undent-region  M-)  Remove a leading indent from a region
universal-argument             ^U  Execute the following command 4 times
unmark-buffer  M-~  Unmark the current buffer (so it is no longer changed)
update-screen  (none)  Force a screen update during macro execution
view-file                      ^X^V  Find a file, and put it in view mode
widen-from-region              ^X>  restores hidden text (see narrow-to-region)
wrap-word (none) Wrap the current word, this is an internal function
write-file                     ^X^W  Write the current buffer under a new file name
write-message (none) Display a string on the command line
yank                           ^Y  yank the kill buffer into the current buffer at the cursor

Command  Binding  Meaning

yank-pop

M-Y Act as yank the first time, but if repeated without moving the cursor the next newest killed text replaces the text just yanked (up to 16 kills saved)

66

MicroEMACS Bindings  MicroEMACS Reference Manual

Appendix D

MicroEMACS Bindings

Below is a complete list of the key bindings used in MicroEMACS. This can be used as a wall chart reference for MicroEMACS commands.

Default Key Bindings for MicroEmacs 3.12

Binding Meaning  Binding Meaning
^A  Move to start of line  ESC A Apropos (list some commands)
^B  Move backward by characters  ESC B Backup by words
^C  Insert space  ESC C Initial capitalize word
^D  Forward delete  ESC D Delete forward word
^E  Goto end of line  ESC E Reset Encryption Key
^F  Move forward by characters  ESC F Advance by words
^G  Abort out of things  ESC G Go to a line
^H  Backward delete
^I  Insert tab/Set tab stops
^J  Insert <NL>, then indent
^K  Kill forward  ESC K Bind Key to function
^L  Refresh the screen  ESC L Lower case word
^M  Insert <NL>  ESC M Add global mode
^N  Move forward by lines  ESC N Goto End paragraph
^O  Open up a blank line
^P  Move backward by lines  ESC P Goto Begining of paragraph
^Q  Insert literal  ESC Q Fill current paragraph
^R  Search backwards  ESC R Search and replace
^S  Search forward  ESC S Suspend (BSD only)
^T  Transpose characters
^U  Repeat command four times  ESC U Upper case word
^V  Move forward by pages  ESC V Move backward by pages
^W  Kill region  ESC W Copy region to kill buffer

MicroEMACS Reference Manual  MicroEMACS Bindings

Binding Meaning  Binding  Meaning
ESC X  Execute named command
^Y  Yank back from kill buffer  ESC Y  Yank pop
^Z  Move backward by pages  ESC Z  Save all buffers and exit
ESC ^C Count words in region  ESC ~  Unmark current buffer
ESC ^E Execute named procedure  ESC ?  Display the file emacs.hlp
ESC ^F Goto matching fence  ESC !  Reposition window
ESC ^H Delete backward word  ESC <  Move to start of buffer
ESC ^K Unbind Key from function  ESC >  Move to end of buffer
ESC ^L Reposition window  ESC .  Set mark
ESC ^M Delete global mode  ESC space Set mark
ESC ^N Rename current buffer  ESC rubout Delete backward word
ESC ^R Search & replace w/query rubout Backward delete
ESC ^S Source command file
ESC ^V Scroll next window down
ESC ^W Delete Paragraph
ESC ^X Execute command line
ESC ^Y Dump the killed text in the ring
ESC ^Z Scroll next window up
^X <  Narrow-to-region  ^X ?  Describe a key
^X >  Widen-from-region  ^X !  Run 1 command in a shell
^X =  Show the cursor position  ^X @  Pipe shell command to buffer
^X ^Enlarge display window  ^X #  Filter buffer thru shell filter
^X 0  Delete current window  ^X $  Execute an external program
^X 1  Delete other windows  ^X (  Begin macro
^X 2  Split current window  ^X )  End macro
^X ^A  Write buffer to the end of file  ^X A  Set variable value
^X ^B  Display buffer list  ^X B  Switch a window to a buffer
^X ^C  Exit MicroEMACS  ^X C  Start a new command processor
^X ^D  Detab line  ^X D  Suspend MicroEMACS (BSD4.2 only)
^X ^E  Entab line  ^X E  Execute macro
^X ^F  Find file  ^X F  Set fill column
^X G  Display variable
^X ^I  Insert file
^X K  Delete buffer
^X ^L  Lower case region
^X ^M Delete Mode  ^X M  Add a mode
^X ^N  Move window down  ^X N  Rename current filename
^X ^O  Delete blank lines  ^X O  Move to the next window
^X ^P  Move window up  ^X P  Move to the previous window
^X ^R  Get a file from disk  ^X R  Incremental reverse search
^X ^S  Save current file  ^X S  Incremental forward search
^X ^T  Trim line  (Incremental search
^X ^U  Upper case region  not always available)
^X ^V  View file

Binding Meaning  Binding Meaning
^X ^W Write a file to disk  ^X W  Resize Window
^X ^X Swap ”.” and mark  ^X X  Use next buffer
^X Y  Move the position in the ring
^X ^Z Shrink window  ^X Z  Enlarge display window

Usable Modes

WRAP  Lines going past right margin ”wrap” to a new line
VIEW  Read-Only mode where no modifications are allowed
CMODE Change behavior of some commands to work better with C
EXACT Exact case matching on search strings
OVER  Overwrite typed characters instead of inserting them
CRYPT Current buffer will be encrypted on write, decrypted on read
MAGIC Use regular expression matching in searches
ASAVE Save the file every 256 inserted characters
SPELL Invoke MicroSPELL to check for spelling errors
REP  Similar to OVER, handles double-byte characters and tabs differently

WHITE/CYAN/MAGENTA/YELLOW/BLUE/RED/GREEN/BLACK
/GREY/GRAY/LRED/LGREEN/LYELLO/LBLUE/LMAGENTA/LCYAN

Sets foreground color

white/cyan/magenta/yellow/blue/red/green/black
/grey/gray/lred/lgreen/lyello/lblue/lmagenta/lcyan

Sets background color 70

Numeric Arguments to Commands  MicroEMACS Reference Manual

Appendix E

Numeric Arguments to Commands

In general, preceding a MicroEMACS command with a numeric argument n causes the command to be executed n times. However, there are a great many commands for which this has no effect, simply because it would make no sense for the command to be executed more than once. There are also commands that take advantage of the numeric arguments to alter their behavior subtly or unsubtly. The following is a list of these commands. Commands that are not affected at all by numeric arguments are listed afterwards.

backward-character A negative argument invokes forward-character. change-screen-size With no arguments, the number of rows defaults to the largest. Otherwise, set the screen size

to n.

change-screen-width With no arguments, the number of columns defaults to the largest. Otherwise, set the screen

width to n.

clear-and-redraw With an argument, centers the window around the current cursor position. cycle-ring With a negative argument, moves forwards in the kill ring, otherwise moves backwards. delete-next-character A negative argument invokes delete-previous-character. delete-next-word With an argument of 0, will not delete the whitespace trailing the deleted word. A negative

argument will cause nothing to happen.

delete-previous-character A negative argument invokes delete-next-character. delete-previous-word An negative or zero argument will cause nothing to happen. detab-region Without an argument, detab-region changes hard tabs to spaces in the lines between the mark

and the cursor. With an argument n, the commands detab n lines - forward if n is positive, backwards if not.

end-of-word A negative argument invokes next-word. entab-region Without an argument, entab-region changes spaces to hard tabs in the lines between the mark

and the cursor. With an argument n, the commands entab n lines - forward if n is positive, backwards if not.

exchange-point-and-mark Swap the current cursor position and mark number n. Without an argument, n defaults to 0. exit-emacs Providing a numeric argument n causes two things to happen. First, no checking for modified

buffers will occur. Second, MicroEMACS exits with a status of n.

MicroEMACS Reference Manual  Numeric Arguments to Commands

forward-character A negative argument invokes backward-character. goto-line An argument n will be taken as the line number to go to. Without an argument, you will be

asked for a line number. In either case, the line number must be 1 or greater.

goto-mark Go to mark number n. Without an argument, n defaults to 0. grow-window A negative argument invokes shrink-window. An argument of 0 causes no action. handle-tab Without an argument, handle-tab deals with the tab character, whether it should be a single

”hard” tab, or expanded as spaces. With an argument n, $softtab is set to n.

hunt-backward The command will hunt n times. The command will report failure if it cannot find its pattern the nth time, even if has found an occurrence of the pattern before number n. A negative argument invokes hunt-forward.

hunt-forward The command will hunt n times. The command will report failure if it cannot find its pattern the nth time, even if has found an occurrence of the pattern before number n. A negative argument invokes hunt-backward.

kill-to-end-of-line With no argument n, the command deletes all characters to the end of the line. If it is already at the end of the line, it will delete the newline. With a positive n as an argument, the command will delete n complete lines, newline character and all, starting from the cursor. With n equal to zero, the command deletes all text from the cursor to the beginning of the line, but will not delete past the newline character. A negative n is illegal.

list-buffers With a numeric argument, INVISIBLE buffers are also listed. move-window-down With a negative argument, invokes move-window-up. move-window-up With a negative argument, invokes move-window-down. next-buffer With an argument n, the nth buffer after the current one is selected, and read in if necessary.

Any buffers in between the current buffer and the target buffer that have not yet been read in are read.

next-line A negative argument invokes previous-line. next-page Without an argument, the window is scrolled forward by a full page. With an argument n, the

window is scrolled forwards by n lines. The cursor is placed on the upper left hand corner. Negative arguments invoke previous-page.

next-paragraph A negative argument invokes previous-paragraph. next-window With a positive argument n, the nth window from the top becomes the working window. With

a negative argument, the nth window from the bottom becomes the working window.

next-word A negative argument invokes previous-word. pop-buffer Without an argument, the buffer is simply displayed in its pop-up screen. With an argument,

the buffer is not only displayed, but also given the attribute INVISIBLE.

previous-line A negative argument invokes next-line. previous-page Without an argument, the window is scrolled backward by a full page. With an argument n,

the window is scrolled backwards by n lines. The cursor is placed on the upper left hand corner. Negative arguments invoke next-page.

previous-paragraph A negative argument invokes next-paragraph. previous-window With a positive argument n, the nth window from the bottom becomes the working window.

With a negative argument, the nth window from the top becomes the working window.

previous-word A negative argument invokes next-word. query-replace-string With a numeric argument, n occurrences of the search string may be replaced, depending

upon the user’s response. The count is based on the number of occurrences found, not the number of positive responses from the user.

quick-exit Saves all modifed buffers, and exits with a status of n.

redraw-display With no argument, or when n is 0, the window is adjusted so that the cursor is in the center. When n is positive, the window is adjusted so that the cursor is on the nth line of the screen. When n is negative, the window is adjusted so that the cursor is on the last line of the window, regardless of the magnitude of n.

remove-mark Remove mark number n. Without an argument, n defaults to 0.

replace-string Will replace n occurrences of the search string with the replacement string. Otherwise, with no argument, all occurrences from the cursor position to the end of file are replaced.

resize-window Requires an argument which must be positive.

scroll-next-down A negative argument invokes scroll-next-up.

scroll-next-up A negative argument invokes scroll-next-down.

search-forward The command will search n times. The command will report failure if it cannot find its pattern the nth time, even if it has found an occurrence of the pattern before number n. A negative argument invokes search-reverse.

search-reverse The command will search n times. The command will report failure if it cannot find its pattern the nth time, even if has found an occurrence of the pattern before number n. A negative argument invokes search-forward.

select-buffer Without an argument, the buffer is simply displayed in the window. With an argument, the buffer is not only displayed, but also given the attribute INVISIBLE.

set If using the set command interactively, preceding the command with a numeric argument then makes it unecessary for the command to ask for the variable’s value (it will still ask for the variable’s name). If used in a command line, then the command

         set <variable name> <number>

is identical to

         <number> set <variable name>

set-fill-column With an argument, the fill column is set to n. The default argument is 1.

set-mark Set mark number n. Without an argument, n defaults to 0.

shrink-window A negative argument invokes grow-window. An argument of 0 causes no action.

split-current-window With n = 1, the new upper window becomes the current window. Any other numeric argument makes the new lower window the current window. With no argument, the current window becomes the new upper or lower window depending upon whether the cursor was in the upper or lower half of the old window.

store-macro Since macroes are numbered, a numeric argument must be provided. These numbered macroes are being phased out in preference for named macroes.

store-procedure If the command is provided a numeric argument, it will assume that store-macro is actually being called.

trim-region Without an argument, trim-region removes spaces and tabs from the end of the lines between the mark and the cursor. With an argument n, the commands trim n lines - forward if n is positive, backwards if not.

E.1 Commands unaffected by numeric arguments.

abort-command add-global-mode add-mode append-file apropos back-from-tag-word begin-macro beginning-of-file beginning-of-line bind-to-key buffer-position case-region-lower case-region-upper change-file-name clear-message-line copy-region count-words cycle-screens delete-blank-lines delete-buffer delete-global-mode delete-mode delete-other-windows delete-screen delete-window describe-bindings describe-functions describe-key describe-variables display end-macro end-of-file end-of-line execute-command-line execute-program fill-paragraph filter-buffer find-file find-screen goto-matching-fence help i-shell incremental-search insert-file kill-region macro-to-key mouse-move-down mouse-move-up mouse-region-down mouse-region-up mouse-resize-screen name-buffer narrow-to-region nop pipe-command print re-tag-word read-file restore-window reverse-incremental-search save-file save-window set-encryption-key shell-command suspend-emacs tag-word transpose-characters unbind-key unmark-buffer update-screen view-file widen-from-region wrap-word write-file write-message

Supported machines  MicroEMACS Reference Manual

Appendix F

Supported machines

The following table lists all the hardware/compilers for which I currently support MicroEMACS. This is not exclusive of all machines which MicroEMACS will run on, but I have either run it myself, or had a first hand report of it running.

Hardware  OS  Compiler  Comments
Hardware  OS  Compiler  Comments
VAX 780  UNIX V5  native
UNIX V7  native
BSD 4.2  native  job control supported
VMS  native  SMG & ANSI support
SUN  SUNOS 3 & 4  native

gcc

NCR Tower  UNIX V5  native
IBM-RT PC  BSD 4.3  native
AIX  native
HP9000  UNIX V5  native
Fortune 32:16  UNIX V7  native
IBM-PC  MSDOS  LATTICE 3  Large CODE/Large DATA
2.0 & 3.2  AZTEC 3.4e Large CODE/Large DATA
TURBO C 2.0 LARGE memory model

MSC 6.0 *MWC 86

SCO XENIX  native
HP150  MSDOS  Lattice 2.15  Function key labels
Turbo C 2.0  for the touch screen

MicroEMACS Reference Manual  Supported machines

Hardware  OS  Compiler  Comments

HP110  MSDOS  Lattice 2.15

Aztec 3.4e Turbo C 2.0

Data General 10  MSDOS  Lattice 2.1  Texas Instruments Professional
MSDOS  Lattice 2.15
Amiga  Intuition  Lattice 3.03

Aztec 3.6

ST520  TOS  Mark Williams C Spawns under MSH
Lattice 3.1  (no shell commands)
Fujitsu FMR  MSDOS  MSC 6.0

series

NEC 9800  MSDOS  Turbo 2.0  Function key support
series  MSC 6.0
HP3000 series  MPE  native

Systems to be supported (IE some code is already written:)

Macintosh  System 7  Lightspeed C

*means that I do not own or have access to the listed compiler and/or machine and must rely upon others to help support it.

Function Keys  MicroEMACS Reference Manual

Appendix G

Function Keys

All environments now support a set of machine independant bindings for function keys. Below is a list of these bindings (not all of these are supported on all systems).

Function keys in MicroEmacs

function Function ^function Alt-function
f1)  FN1  S-FN1  FN^1  A-FN1
f2)  FN2  S-FN2  FN^2  A-FN2
f3)  FN3  S-FN3  FN^3  A-FN3
f4)  FN4  S-FN4  FN^4  A-FN4
f5)  FN5  S-FN5  FN^5  A-FN5
f6)  FN6  S-FN6  FN^6  A-FN6
f7)  FN7  S-FN7  FN^7  A-FN7
f8)  FN8  S-FN8  FN^8  A-FN8
f9)  FN9  S-FN9  FN^9  A-FN9
f10)  FN0  S-FN0  FN^0  A-FN0
home) FN<  FN^<
CsUp) FNP  FN^P
PgUp) FNZ  FN^Z
CsLf) FNB  FN^B

5 )

CsRt) FNF  FN^F
End) FN>  FN^>
CsDn) FNN  FN^N
PgDn) FNV  FN^V
Ins)  FNC  FN^C
Del)  FND  FN^D

MicroEMACS Reference Manual  Function Keys

78

Machine Dependent Notes  MicroEMACS Reference Manual

Appendix H

Machine Dependent Notes

This appendix lists some notes specific to individual implementations of MicroEMACS. Every attempt has been made to allow EMACS to be identical on all machines, but we have also tried to take advantage of function keys, cursor keys, mice, and special screen modes where possible.

H.1 IBM-PC/XT/AT and its clones

The IBM-PC family of computers is supported with a variety of different display adapters. EMACS will attempt to discover what adapter is connected and use the proper driver for it. Below is a list of the currently supported video adapters:

Adapter  $sres  Original mode used
Monochrome Graphics Adapter MONO MONO
Color Graphics Adapter CGA CGA
CGA40 CGA40
Enhanced Graphics Adapter  EGA  CGA
Video Graphics Adapter  VGA  CGA

VGA12

If a driver for a Microsoft compatable mouse is installed on the system, EMACS will use the mouse in text mode and allow the user all the standard mouse functions. The mouse cursor will appear to be a block of color in the color opposite of it’s background.

EMACS also takes advantage of various function keys and the keys on the keypad on an IBM-PC. The function keys are initially not bound to any particular functions (except by the emacs.rc startup file), but the keypad keys do default to the following:

Keypad key  Function
Home  beginning-of-file
CSRS UP  previous-line
Pg Up  previous-page
CSRS LEFT backward-character
CSRS RIGHT forward-character
End  end-of-file
CSRS DOWN next-line
Pg Dn  Next-page

All these special keys are indicated in EMACS macroes by use of the FN prefix. Below is a list of many of the keys and the codes used to specify them. Also the codes may be gotten by using the describe-key (^X ?) command on the suspect key.

Compiling under TURBO C

To compile MicroEMACS under TURBO C, set the TURBO integrated environment with the following options:

MicroEMACS Reference Manual  Machine Dependent Notes

Memory model  LARGE
Floating point  NONE
Default char type  UNSIGNED
Data alignment  BYTE
Merge duplicate strings ON Standard stack frame off
Test stack overflow  off
Optimize for  SIZE
Use register optimization  ON
Register optimization  ON
Jump optimization  ON
Initialize segments  OFF
Stack warnings  OFF

Names: Code names

Segment name  *

H.2 HP 150

This machine from Hewlett Packard is very unusual for an MSDOS machine. It has a touch screen and is very function key oriented. An additional command, label-function-key allows you to place labels on the on screen function key labels. A numeric argument indicates which function key to label (one through eight) and then the program prompts for a 16 character label, which will be used as two lines of eight characters. To label function key three with ”save file” from a macro, you would use:

     3 label-function-key "save   file"

Notice the 4 spaces after ”save”. This forces ”file” to begin on the second line of the label.

H.3 Atari 520/1040ST

The ATARI ST family of computers have a dual personality. They may use either a monochrome or a color screen. EMACS supports two screen resolutions on each monitor.

NOTE

When you set MicroEMACS up on your system, please remember to install it on the desktop as a GEM application. If you have EMACS set as a TOS application, the mouse will not function properly, and EMACS will alert you to this problem by beeping the bell.

Monitor $sres  size  #color $palette format

 Color  LOW  40x25 16  000111222333444555666777
          MEDIUM 80x25 4  000111222333

Mono  HIGH  80x25 2  000

          DENSE  80x50 2  000

The $palette environment variable can be used to change what color is associated with each color name. With a color monitor, each group of three digits indicates an octal number specifying the RED, GREEN and BLUE levels of that color. Each color digit can vary from 0 to 7. For example, the initial setting of $palette in LOW resolution is:

              000700070770007707077777

which broken up is:

              000 700 070 770 007 707 077 777

which means:

              000      Black
              700      Red
              070      Green
              770      Yellow
              007      Blue
              707      Magenta
              077      Cyan
              777      White

Also the mouse buttons are bound to mouse functions as described in the chapter about mice. The cursor keys and the function keys are bound similarly to IBM-PC.

Files generated by EMACS on the ATARI ST have a single return character at the end of each line, unlike the desktop files which want to have two returns. This makes it display files strangely from GEM’s [SHOW] option, but makes the files port to other computers much nicer. When compiling MicroEMACS, the ADDCR symbol in estruct.h will cause emacs to generate line ending sequences compatible with GEM.

Currently, when operating under the Mark Williams MSH program, EMACS can shell out and perform external commands. This capability will be added later for the Beckmeyer shell and under GEMDOS.

H.4 Amiga 1000

The Commodore AMIGA 1000 version of MicroEMACS does fully support the mouse, window resizing and the close gadget. It runs in medium resolution, using the colors defined for the workbench.

Note about Compiling MicroEMACS

If you are compiling the sources on the AMIGA to produce an executable image, and you are using the Lattice compiler, be sure to give the CLI command STACK 40000 before compiling to make sure the compiler has sufficient stack space to successfully complete compilation.

H.5 UNIX V5, V7, and BSD4.[23]

MicroEMACS under UNIX utilizes the TERMCAP library to provide machine independent screen functions. Make sure that termcap is available and properly set on your account before attempting to use MicroEMACS.

Under systems which support job control, you can use the ^XD suspend-emacs command to place EMACS into the background. This carries a much smaller overhead than bringing up a new shell under EMACS. EMACS will properly redraw the screen when you bring it back to the foreground.

If the symbol VT100 has been set to 1 in the estruct.h options file, EMACS will recognize the key sequence <ESC>[ as the lead in sequence for the FN function key prefix.

With the addition of some very machine/operating system specific code, EMACS can prevent two or more people from modifying the same file at the same time. The upper level of a set of functions to provide file locking exists in the source file lock.c. It requires two machine specific functions written and linked into EMACS for it to operate properly.

              char *dolock(fname)
              char *fname;
              dolock() locks a file, preventing others from modifying it. If
              it succeeds, it returns NULL, otherwise it returns a pointer to
              a string in the form "LOCK ERROR: explanation".
              char *undolock(fname)
              char *fname;
              undolock() unlocks a file, allowing others to modifying it. If
              it succeeds, it returns NULL, otherwise it returns a pointer to
              a string in the form "LOCK ERROR: explanation".

H.6 DEC VMS operating system

TERMINALS

Depending upon the options set in ESTRUCT.H, MicroEMACS uses either the capabilities of VMS SMG, working with any terminal that is defined in SMGTERMS.TXT or TERMTABLE.TXT (see your SMG manual for more information), or the ANSI escape sequences. Full keyboard support, including function keys, is provided for VT100 and VT200 series compatible terminals. Mouse support is provided under the ANSI version only at this time. Mouse support is provided for the VSII workstation’s VT220 terminal emulator, and other terminal emulators that use the same escape sequences for mouse control. (There is some partial support for the BBN BitGraph mouse sequences in the sources, but this is not yet complete). Terminals may have up to 100 lines and 160 columns.

The maximum terminal size is 256 columns and 72 rows. If you run MicroEMACS on a terminal that is larger than this, MicroEMACS will reduce it to these limits while you are editing.

Flow control

Some terminals will require the use of XON/XOFF flow control when used with MicroEMACS. When XON/XOFF flow control is used, you will not be able to use functions bound to ^S or ^Q, and should use bind-to-key to put these functions on other keys. MicroEMACS does not change the flow control characteristics of your terminal line while it is running. If your terminal requires flow control, you should:

     $ SET TERM/HOSTSYNC/TTSYNC

before entering MicroEMACS. If you are on a VSII emulated workstation terminal, are using the SSU multi-session protocol (VT330 and VT340 with SSU enabled), or are certain that your terminal does not require XON/XOFF flow control, you should

     $ SET TERM /HOSTSYNC/NOTTSYNC

This will allow you to use ^S and ^Q for MicroEMACS commands. Note that if you are using a VSII with VWS V3.2 or later, you must leave the /HOSTSYNC enabled in order for the cross/session cut and paste capability to work properly.

KEYBOARD

The VMS version understands the LK201 functions of VT200 series, VT300 series, and compatible terminals and terminal emulators, and allows you to bind to them as function keys. In addition, the VT100 numeric keypad, in application mode, is available as function keys. MicroEMACS will only put the keypad into application mode for you if the KEYPAD option is set in ESTRUCT.H. In this situation, MicroEmacs will detect your kepad’s state, and restore it to that state upon exiting. If MicroEMACS has not been compiled with this option, you may still put the keypad into application mode by issuing the command SET TERM /APPLICATION before entering MicroEMACS.

VT200 keys

Note that F1 through F5 are local function keys on DEC terminals.

F6  = FN6  FIND = FNS
FN7  = FN7  INSERT = FNC
F8  = FN8  REMOVE = FND
F9  = FN9  SELECT = FN@
F10  = FN0  PREV = FNZ
F11  = S-FN1 NEXT = FNV
F12  = S-FN2 Arrow Up = FNP
F13  = S-FN3 Arrow Down = FNN
F14  = S-FN4 Arrow Right = FNF
HELP (F15) = S-FN5 Arrow Left = FNB
DO (F16)  = S-FN6
F17  = S-FN7
F18  = S-FN8
F19  = S-FN9
F20  = S-FN0

VT100 and VT200 numeric keypad in application mode

PF1 = FN^1 PF2 = FN^2 PF3 = FN^3 PF4  = FN^4
7  = A-7  8  = A-8  9  = A-9  -  = A–
4  = A-4  5  = A-5  6  = A-6  ,  = A-,
1  = A-1  2  = A-2  3  = A-3  ENTER = A-E
0  = A-0  .  = A-.

WARNING

The VMS version contains code for interpreting function keys that are sent as Ansi sequences that begin with the ESC character. Because of this, MicroEMACS cannot process an incoming ESC until it knows what character follows it. This can cause problems with terminating search and replace strings. If you use ESC as the meta-prefix character (which is the default) you must type one additional keystroke following ESC before emacs will recognize that you have edited the search command prompt, and are continuing. (The additional character is processed normally be MicroEMACS, it is NOT discarded.) MicroEMACS must wait long enough for the network delay that might be involved between seeing the ESC and seeing the characters that follow it. If holding down one of the arrow keys causes characters to drop into your file, then you may want to alter the delay yourself. The logical variable MICROEMACS$SHORTWAIT may be set to vary that delay. The default delay is 400ms (4 tenths of a second). The equivalent value in MICROEMACS$SHORTWAIT is 4000000.

Special case for BBN BItGraph

If you are using the BBN BitGraph, execute the following commands before entering MicroEMACS, and you will get mouse support:

             \$ esc[0,8] = 27
             \$ microemacs\$mouse\_enable == esc+":5;6;L"+esc+":0;63;;;;;;;;;9;16;c"
             \$ microemacs\$mouse\_disable == esc+":5;1;L"+esc+":0;0c"
             \$ exit

Do NOT do this for any other terminals.

Search List for EMACS.RC

VMS MicroEMACS will first search logical name MICROEMACS$LIB:, then SYS$LOGIN:, then the current directory, and finally sys$sysdevice:[vmstools] when looking for startup files or help files.

Please use MICROEMACS$LIB:, and allow the secondary search of [vmstools] to become archaic. If desired, MICROEMACS$LIB may be defined to be a VMS search list that first searches a user directory, and then a system directory.

Generally, you should create a private directory where you keep all your .CMD files, and in your LOGIN.COM $DEFINE a logical name to point to this area.

In addition to whatever commands you have in your EMACS.RC file, one command you should certainly include is set $ssave FALSE. The ”safe save” mechanism, which writes a buffer to a temporary file, deletes the old version of a file, and then moves the temporary file to its permanent name, works wonderfully on most systems, but makes no sense on VMS, which maintains older versions of a file.

Using MicroEMACS as a subprocess

MicroEmacs can now be kept in a subprocess. You can arrange to start emacs only once in a job, and to re-attach to it each time you want to use it. This is optional. To use this feature, install MicroEMACS in the following way:

1. MicroEMACS contains two images. ME.EXE is a small program for starting and stopping the Emacs subprocess. The source for ME. is in ME.C, and should not be linked into MESHR.EXE. MESHR.EXE is the actual MicroEMACS image. The name MESHR is required for MAIL/NOTES support, see next section for details.

2. Make sure that the SYS$SHARE search list includes MESHR.EXE. If you don’t have the privilages to move MESHR.EXE into SYS$SHARE, you can $ DEFINE the MESHR logical name to be the full name and location of the MESHR.EXE program. For example, you could store all of these programs in the MICROEMACS$LIB: search list, and say:

     $ DEFINE MESHR microemacs$lib:meshr.exe

3. Put ME.EXE in MICROEMACS$LIB and the following line in your LOGIN.COM:

     $ me :== $microemacs$lib:me

4. Put a line in your EMACS.RC that will

     bind-to-key suspend-emacs ^C ; use your usual exit-emacs key

Now, use the $ ME command to invoke microemacs. Subseqeuent invocations in the same job will re-use the existing subprocess. You can use the full capabilty of the microemacs command line in the first and in all subsequent invocations of ME.

WARNING

MicroEMACS will ALWAYS read in new copies of any files you specify on the command line, even if you are already editing it. If you edit a file a second time with the same MicroEMACS, you will get a NEW buffer with ANOTHER copy of the file. The old buffer is still there also. It is easy, in this situation, to accidently edit in a WRONG BUFFER, and if you write out an obsolete buffer, you will lose earlier edits! This is considered a bug and may be fixed in a later version of MicroEMACS. To avoid this situation, do not specify a file on the command line if MicroEMACS already has that file in a buffer. Use the find-file MicroEMACS command instead.

Using MICROEMACS with MAIL and NOTES:

With VMS V5 and later versions, the MAIL interface to Microemacs is much simplified. With VMS V5, the MESHR.EXE image does NOT have to be installed as a known image to be used as a callable editor from MAIL. Therefore, to use MicroEMACS as your VMS MAIL editor, simply add the following lines to your LOGIN.COM:

                       \$ DEFINE MAIL\$EDIT CALLABLE\_ME
                       \$ MAIL :== MAIL/EDIT

and make sure that the SYS$SHARE search list includes MESHR.EXE. If you don’t have privs or permission to move MESHR.EXE into SYS$SHARE, you can $ DEFINE the MESHR logical name to be the full name and location of the MESHR.EXE program. For example, you could store all of these programs in the MICROEMACS$LIB: search list, and say:

     $ DEFINE MESHR microemacs$lib:meshr.exe

Note that this is the same location as is required for using kept MicroEMACS.

To abort sending a message, exit MicroEMACS without writing out the mail message file.

To use MicroEMACS as your VAX NOTES editor, issue the following command to VAX NOTES:

     NOTES> SET PROFILE/EDIT=(ME,CALL)

Note, if you are still in the dark ages of VMS V4, you will have to either install MESHR as a known image, or following the original ”Second way” instructions given in the existing appendix F.6 of the older MicroEMACS manual (previous to version 3.10).

Second way, as described in older versions

In the event that you cannot get your system manager to INSTALL MicroEMACS as known image, you can use the following technique:

1. In MICROEMACS$LIB:MEMAIL.COM, put the following command file:

$! Use on VAX/VMS as MAIL$EDIT for using MicroEMACS as mail editor. $ if ””P1’” .NES. ” NL:” then if ””P1’” .NES. ”” then copy ’P1’ ’P2’ $ define/user sys$input sys$output $ me ’P2’ $ exit

This file may have come with your MicroEMACS kit.

2. In your LOGIN.COM, put the following lines:

         \$       me :== \$MICROEMACS\$LIB:MESHR.EXE ! Assumes meshr.exe is there
         \$       define mail\$edit microemacs\$lib:me\_edit.com

3. In NOTES, give the command

         NOTES> SET PROFILE/EDIT=(@MicroEMACS$lib:me edit.com,SPAWN)

System messages and EMACS

MicroEMACS will intercept system broadcast messages and display them on the message line after any input from the user. These message are stored in an INVISIBLE buffer named [-messages-]. To view these at your convenience, use the following procedure:

     ;
     ; Show any system messages MicroEMACS may have intercepted.
     ; The numeric prefix of pop-buffer ensures its invisibility.
     ;
     store-procedure "Show\_messages"
              1 pop-buffer "[-messages-]"
     !endm

Building MicroEMACS for VMS

The configuration options are set in file estruct.h:

• Under the category of ”Machine/OS definitions”, set VMS to ”1” and all others to ”0”.

• Under ”Compiler definitions”, set all selections to ”0”. Selecting VMS implies that you are using VAXC.

• Under ”Special keyboard definitions”, be sure ”VT100” is set to ”0”. This option is not required for the VMS version, it is for other systems using ANSI terminal support. VMS in combination with SMG or ANSI already handles the special characteristics of Ansi keyboards.

• Under ”Terminal Output definitions”, set either ANSI or SMG to ”1” and all others to ”0”. As stated previously, only ANSI supports the mouse at this time.

• Under ”Configuration options”, you may select as you wish, with the following notes:

COLOR support does not exist for VMS, even when using color workstations.

MOUSE support should be enabled if you have any VSII workstations. Only supported under the ANSI driver.

KEYPAD support recognises whether your keypad is already in application mode or not, and puts your keypad in its correct state on exit.

XONDATA automatically allows you to use control-S or control-Q in MicroEMACS, by disabling the TTSYNC characteristic. This option should not be set if MicroEMACS might be used on DecStations or VT100s. It also should not be used with slow terminals or terminal emulators connected to fast terminal lines.

RMSIO support should absolutely be used. This option allows the writing and reading of files in VMS’s variable-length format, as opposed to STREAM-LF, and cuts down on file writing and reading time by approximately two thirds.

OPTMEM support may be used on VMS versions 5.0 and higher. It substitutes the C library’s memory allocation calls for the native VAX calls, and gives a speed improvement.

If you have MMS, you can use the supplied DESCRIP.MMS to build MicroEMACS. Otherwise, the command file MEMAKE.COM has been provided. These files assume that you are using SMG as your terminal driver. If you are using ANSI, then you must replace SMG with ANSI in the command and opt files. If you do not have MMS or are missing MEMAKE.COM, simply compile each module with ”CC”, and link with the command:

     $ LINK MESHR/OPTION/SHARE

Note that the executable filename must end in ”SHR” in order for MicroEMACS to be used as a callable editor from MAIL or NOTES (Method 1 above).

If you edit any of the Emacs sources, note that any global or external data must be declared as ”noshare” in order for the VMS callable editor support to work properly. This applies to all global data used in the VMS version, but not to routines or to ”static” data. The ”noshare” declaration is #define’d away on non-VMS systems. If you fail to do this, VMS will not allow you to INSTALL MicroEMACS as a sharable library.

90

Mode Flags  MicroEMACS Reference Manual

Appendix I

Mode Flags

The two environment variables, $cmode and $gmode, contain a number the corresponds to the modes set for the current buffer and the editor as a whole. These are encoded as the sum of the following numbers for each of the possible modes:

WRAP      1 Word wrap
CMODE     2 C indentation and fence match
SPELL     4 Interactive spell checking (Not Implemented Yet)
EXACT     8 Exact matching for searches
VIEW     16 Read-only buffer
OVER     32 Overwrite mode
MAGIC    64 Regular expressions in search
CRYPT   128 Encryption mode active
ASAVE   256 Auto-save mode
REP     512 Similar to OVER

So, if you wished to set the current buffer to have CMODE, EXACT, and MAGIC on, and all the others off, you would add up the values for those three, CMODE 2 + EXACT 8 + MAGIC 64 = 74, and use a statement like:

    set $cmode 74

or, use the binary or operator to combine the different modes:

    set $cmode &bor &bor 2 8 64

Internal Flags

Some of the ways EMACS controls its internal functions can be modified by the value in the $gflags environment variable. Each bit in this variable will be used to control a different function. GFFLAG 1 If this bit is set to zero, EMACS will not automatically switch to the buffer of the first file after

executing the startup macroes.

GFSDRAW   2 If this bit is set to one, supress redraw events.

Current buffer flags

The $cbflags environment variable allows the user to modify some of the characteristics of the current buffer. The various characteristics are encoded as the sum of the following numbers:

MicroEMACS Reference Manual  Mode Flags

BFINVS    1 Internal invisible buffer
BFCHG     2 Changed since last write
BFTRUNC   4 buffer was truncated when read
BFNAROW   8 buffer has been narrowed

Only the invisible and changed flags can be modified by setting the $cbflags variable. The truncated file and narrowed flags are read only.